### v1

# sample.bib

Abstract.
We consider the convergence of ESD for non-Hermitian random band matrices with independent entries to the circular law, where the bandwidth scales like $n^{\gamma}$ and $n$ is the matrix size. We prove that the circular law limit holds provided that $\gamma>\frac{5}{6}$ for a very general class of inhomogeneous matrix models with Gaussian distribution and doubly stochastic variance profile, and provided that $\gamma>\frac{8}{9}$ if entries have symmetric subGaussian distribution. This improves previous works which essentially require $\gamma>\frac{32}{33}$. We also prove an extended form of product circular law with a growing number of matrices. Weak delocalization estimates on eigenvectors are also derived. The new technical input is new polynomial lower bounds on some intermediate small singular values.
1. Introduction
For an $n\times n$ matrix $A$ with eigenvalues $\lambda_{1}(A),\cdots,\lambda_{n}(A)\in\mathbb{C}$ we denote its empirical measure of eigenvalues by
$\mu_{A}:=\frac{1}{n}\sum_{i=1}^{n}\delta_{\lambda_{i}(A)},$
where $\delta_{z}$ is the Dirac measure at point $z$.
For a symmetric random matrix $A$, the convergence of $\mu_{A}$ to the semicircle distribution under mild assumptions dates back to the celebrated work of Wigner [wigner1958distribution]. The spectrum of symmetric random matrices can essentially be determined via the method of moments. For non-Hermitian random matrix $A$, the convergence of $\mu_{A}$ does not follow from method of moments arguments (see for example [MR2908617]), and fundamentally new ideas are needed. When all the entries of $A=(a_{ij})_{1\leq i,j\leq n}$ are i.i.d. with mean zero and variance one, the circular law theorem, i.e. the convergence of $\mu_{A}$ to the uniform distribution on the unit disk in the complex plane, is one of the fundamental theorems in modern random matrix theory. This circular law was proven under increasing generality in a series of works including [girkoarticle], [MR1428519], [tao2008random], [WOS:000281425000010], and we refer to the references in these works and [MR2908617] for a more complete list. Beyond dense matrices with i.i.d. entries, the circular law was subsequently proven for sparse i.i.d. matrices [MR3980923], [MR3945840] [sah2023limiting], for sparse regular digraphs [MR4195739], and for some random matrices with inhomogeneous variance profile [MR3878135]. A more complete list of references can be found in the review [MR4680362].
Despite all the recent efforts in proving the circular law, effective techniques for structured non-Hermitian random matrices with inhomogeneous variance profile are still out of reach. In these models, a large part of entries are always set to be zero. (Note that for dense inhomogeneous matrices [MR3857860], [MR3878135] developed a general technique to prove circular law, but the approach becomes ineffective for sparse inhomogeneous models). The guiding examples of sparse inhomogeneous matrices are random band matrices, which we will define shortly. Their Hermitian analogue, namely Hermitian Random band matrices, play a fundamental role in modern mathematical physics. They serve as prototypical models of Anderson transition: when the bandwidth scales like $n^{\gamma}$ for $\gamma>\frac{1}{2}$, the eigenvectors are expected to be delocalized and eigenvalues have GOE local statistics; whereas a localized phase and Poisson statistics should occur when $\gamma<\frac{1}{2}$. A list of recent mathematical progress in justifying these conjectures can be found in reviews [MR3966510] and [MR3966868], and also the more recent work [MR4736267].
In this work we study non-Hermitian versions of random band matrix, where even global universality, i.e. the convergence of the empirical spectral density to the circular law, remains largely unproven. In recent works [jain2021circular] [tikhomirov2023pseudospectrum] the authors established circular law for certain non-Hermitian band matrices with bandwidth $n^{\gamma}$ with $\gamma$ sufficiently close to 1: more precisely they require $\gamma>\frac{32}{33}$. The main contribution of this paper is to consider much narrower bandwidth and expand the variety of models where global universality holds: we show $\gamma>\frac{5}{6}$ is enough for the circular law for Gaussian entries, and for a very broad class of inhomogeneous matrix models with doubly stochastic variance profile.
1.1. Circular law for explicit models
We introduce a family of specific band matrix models.
Definition 1.1.
(Block band matrix) Let $X$ be an $n\times n$ matrix. We say $X$ is sampled from a block band matrix ensemble if it has the form
$X=\begin{pmatrix}D_{1}&U_{2}&&&T_{m}\\ T_{1}&D_{2}&U_{3}&&\\ &T_{2}&D_{3}&\ddots&\\ &&\ddots&\ddots&U_{m}\\ U_{1}&&&T_{m-1}&D_{m}\end{pmatrix}$ (1.1)
where the unfilled sites are set zero.
The blocks $D_{1},U_{1},T_{1},\cdots,D_{m},U_{m},T_{m}$ are $b_{n}\times b_{n}$ independent random matrices with i.i.d. entries having the same distribution $\frac{1}{\sqrt{3b_{n}}}\xi$, where $\xi$ is a random variable of mean zero and variance one. Also set $m:=n/b_{n}$, where we assume that $n/b_{n}$ is an integer. This normalization ensures that $\mathbb{E}[XX^{*}]=\mathbb{E}[X^{*}X]=\mathbf{1}$.
Definition 1.2.
(Periodic band matrix) Let $X$ be an $n\times n$ square matrix. We say $X$ is sampled from a periodic band matrix model with bandwidth $d_{n}$ if $x_{ij}=0$ for any $\frac{d_{n}-1}{2}<|i-j|<n-\frac{d_{n}-1}{2}$, and $x_{ij}$ are i.i.d. copies of $\frac{1}{\sqrt{d_{n}}}\xi$ if $|i-j|\leq\frac{d_{n}-1}{2}$ or $|i-j|\geq n-\frac{d_{n}-1}{2}$. Here $\xi$ is a random variable of mean zero and variance one. We implicitly assume $\frac{d_{n}-1}{2}$ is an integer.
Definition 1.3.
A real or complex random variable $\xi$ of mean zero, variance one is said to be $K$-subGaussian if
$\mathbb{E}\exp(|\xi|^{2}/K^{2})\leq 2.$
Throughout the paper we make a further technical assumption on $\xi$ that significantly simplifies the computations after a truncation procedure:
Assumption 1.4.
The random variable $\xi$ satisfies the following condition: there exists a sufficiently large $D_{0}(\xi)>0$ such that, for any $D>D_{0}(\xi)$,
$\mathbb{E}[\xi 1_{|\xi|\leq D}]=0.$ (1.2)
Typical examples covered by Assumption 1.4 are all a.s. bounded random variables (such as Bernoulli random variable and uniform measure on a compact interval), and all random variables $\xi$ with a symmetric distribution $\xi\equiv^{law}-\xi$. This assumption is possibly not necessary, but is made so that we have a much tighter concentration inequality. Condition (1.2) is weaker than saying $\xi$ has a symmetric distribution, but merely that (1.2) has a symmetric law at infinity.
The main results of this paper are as follows.
Theorem 1.5.
Let $X$ be sampled from the block band matrix model in Definition 1.1. Assume that either one of the following is true
  1. (1)
    $\xi$ is a real or complex Gaussian variable and $b_{n}\geq n^{\frac{5}{6}+\epsilon}$ for some $\epsilon>0$, or
  2. (2)
    $\xi$ has all $p$-th moment finite for all $p\in\mathbb{N}_{+}$, satisfies (1.2) and $b_{n}\geq n^{\frac{8}{9}+\epsilon}$ for some $\epsilon>0$,
Then $\mu_{X}$ converges in probability to the uniform measure on the unit disk in the complex plane.
This theorem was previously proven in [jain2021circular] for $b_{n}\gg n^{\frac{32}{33}}$ but with a relaxed moment condition on $\xi$ (assuming $4+\epsilon$ finite moments).
No density assumption on $\xi$ is imposed here: in particular, in this theorem we can take $\xi$ to be the Bernoulli measure supported on only two points.
For the periodic band matrix model (1.2), we prove a similar result where we additionally require that (i) $\xi$ has a bounded density and (ii) $\xi$ has subGaussian tails. These two constraints arise from the least singular value estimate in [tikhomirov2023pseudospectrum].
Theorem 1.6.
Let $X$ be sampled from the periodic band matrix model in Definition 1.2. Assume that either one of the following is true: for some $\rho_{0}>0$,
  1. (1)
    $\xi$ is a real or complex Gaussian variable and $b_{n}\geq n^{\frac{5}{6}+\epsilon}$ for some $\epsilon>0$, or
  2. (2)
    $\xi$ is a $K$-subGaussian for some $K>0$, satisfies (1.2), and $b_{n}\geq n^{\frac{8}{9}+\epsilon}$ for some $\epsilon>0$. Moreover, $\xi$ is either real with a density bounded by $\rho_{0}$, or the real and imaginary parts of $\xi$ are independent and both has a density bounded by $\rho_{0}$.
Then $\mu_{X}$ converges in probability to the uniform measure on the unit disk of the complex plane.
This theorem was previously proven in [tikhomirov2023pseudospectrum] for $b_{n}\gg n^{\frac{33}{34}}$.
We also prove a circular law for the product of a growing number of i.i.d. matrices.
Theorem 1.7.
Let $X_{1},\cdots,X_{m_{n}}$ be independent $n\times n$ random matrices with i.i.d. entries having distribution $\xi$. Let $m=m_{n}$ be $n$-dependent and such that $m_{n}\leq n^{\frac{1}{8}-\epsilon}$ for some $\epsilon>0$. Assume that $\mathbb{E}[\xi]=0,\mathbb{E}[|\xi|^{2}]=1$, $\xi$ satisfies (1.2), and $\xi$ has finite $p$-th moment for all $p\in\mathbb{N}_{+}$. Denote by
$\mathcal{M}:=(X_{1}\cdots X_{m_{n}})^{\frac{1}{m_{n}}}.$
Then in the limit $n\to\infty$, we have that the empirical measure $\mu_{\mathcal{M}}$ converges in probability to the uniform measure on the unit disk of the complex plane.
Remark 1.8.
When $m_{n}$ is fixed and independent of $n$, the circular law of $\mathcal{M}$ has been proved in [MR2861673] under a $2+\eta$-th moment condition. That is, the eigenvalues of $X_{1}\cdots X_{m_{n}}$ converges to the so-called $m_{n}$-fold product circular law. A local circular law has also been proven for product matrices [MR3622892], and subsequently universality of linear statistics for i.i.d matrix products are derived [MR4112718]. Via adapting the proof in [jain2021circular], we can possibly relax the moment condition on $\xi$ in Theorem 1.7 to, say, assuming only a finite six-th moment, at the cost of allowing a much smaller rate of growth of $m_{n}$. As the focus here is on large $m_{n}$, we sacrifice with the moment assumptions.
Optimally, one expects to prove the circular law for all $m_{n}\leq n^{1-\epsilon}$, but this is beyond the techniques in this work. The threshold $m_{n}\sim n$ has reached a lot of recent attention. Informally, $m_{n}\ll n$ is the free probability regime where universality is expected to hold for the product matrix; $m_{n}\sim n$ is the critical transition regime and $m_{n}\gg n$ is the ergodic regime (and there should be no universality). We refer to [MR4401507], [MR4421171] and [MR4580535] for integrable probability perspectives on product random matrices with simultaneously growing $m_{n}$ and $n$, and to [MR4268303] for an ergodic theory perspective.
1.2. Weak delocalization estimates
A major component of our proof is to upper bound the Green function $G(z)$ of a dilated version (3.1) of $X$, at some scales $\Im z\ll 1$. We cannot reach the optimal scale $\Im z\sim n^{-1+\epsilon}$ and are indeed very far from that, but such a bound appear to be new and implies interesting delocalization estimates.
We introduce the most general model where our delocalization statement is valid.
Definition 1.9 (General model).
$X=(x_{ij})$ is a $n\times n$ random matrix with independent entries. For each $(i,j)\in[n]^{2}$ $x_{ij}$ is distributed as $b_{ij}\xi$ where $\xi$ has mean zero, variance one, finite $p$-th moment for all $p\in\mathbb{N}_{+}$ and satisfies (1.2). Set
$b_{n}^{-1}=\sup_{(i,j)\in[n]^{2}}b_{ij}^{2}.$
We assume that $X$ has a doubly stochastic variance profile:
$\mathbb{E}[XX^{*}]=\mathbb{E}[X^{*}X]=\mathbf{1}.$
A special case of this definition is the following: consider $G_{n}=([n],\mathcal{E})$ a $b_{n}$- regular directed graph on $n$ vertices. Then we set $b_{ij}=\frac{1}{\sqrt{b_{n}}}$ if $(i,j)\in\mathcal{E}$ and $b_{ij}=0$ otherwise. All preceding examples in this paper are already covered at this level of generality.
Theorem 1.10 (Weak delocalization).
Consider a random matrix as in Definition (1.9). Assume $b_{n}\geq n^{\epsilon}$ for some sufficiently small $\epsilon>0$. Then for any (sufficiently small) $\mathbf{c}>0$ there holds
  1. (1)
    If $\xi$ has (real or complex) Gaussian distribution. then with probability at least $1-n^{-10}$, all eigenvectors $\psi$ of $X$ with unit $L^{2}$ norm satisfy
    $\|\psi\|_{\infty}\leq(b_{n})^{-1/10}n^{\mathbf{c}}.$
  2. (2)
    For general $\xi$ satisfying assumptions in Definition 1.9, then with probability at least $1-n^{-10}$, all eigenvectors $\psi$ of $X$ with unit $L^{2}$ norm satisfy
    $\|\psi\|_{\infty}\leq(b_{n})^{-1/16}n^{\mathbf{c}}.$
[…]
1.3. Circular law for generic models
Finally we prove that the circular law continues to hold for the very general model in Definition 1.9, i.e. independent of the block structure in Definition 1.1, 1.2.
Theorem 1.11 (Circular law for general model).
Let $X$ be defined as in Definition 1.9, $\xi$ has Gaussian distribution, and $b_{n}\geq n^{\frac{5}{6}+\epsilon}$ for some $\epsilon>0$. Then the circular law holds for $\mu_{X}$.
In this theorem $X$ is only assumed to have a doubly stochastic variance profile: no block or band structure is assumed and the variance of each nonzero entry of $X$ is not assumed identical. To our best knowledge, circular law has not been proven in such a wide generality (but see [MR3878135] and [cook2article] for limiting ESDs of inhomogeneous random matrices with variance upper bounded by $Cn^{-1/2}$, so there are roughly $cn$ nonzero elements per row and column).
The new technical input for this theorem is a convergence rate estimate of Stieltjes transform for the very general model in 1.9. This convergence rate is derived in Theorem 5.1 under more general assumptions on $b_{n}$.
Remark 1.12.
Although Theorem 1.11 is stated for Gaussian entries, its adaptation to the general entry case is straightforward, as in previous theorems. To do this, one can easily formulate a version of Theorem 5.1 for entries with all $p$-th moment finite (using the machinery from [brailovskaya2022universality] instead of [bandeira2023matrix]), and we also need to assume $\xi$ has a bounded density as we use the singular value lower bound in [tikhomirov2023pseudospectrum]. Proposition 3.3 remains valid, and finally we can prove circular law for $K$-subGaussian entries satisfying (1.4) and having a bounded density, under the assumption that $b_{n}\geq n^{\frac{8}{9}+\epsilon}$. The details are omitted.
1.4. Main ideas and outline
We follow the classical Hermitization trick of Girko to establish the circular law, which we refer to [MR2908617] for a through introduction and historical review.
For an $n\times n$ matrix $A$ we denote by $\sigma_{1}(A)\geq\sigma_{2}(A)\cdots\geq\sigma_{n}(A)$ the list of singular values of $A$, arranged in decreasing order.
Thanks to the replacement principle by Tao and Vu (see Theorem 4.1), the main technical step to prove circular law is to show convergence of the logarithmic sums of singular values of shifted matrix $X_{z}:=X-zI_{n}$, $\sum_{i=1}^{n}\log\sigma_{i}(X-zI_{n})$, towards $\sum_{i=1}^{n}\log\sigma_{i}(G-zI_{n})$ for a.e. $z\in\mathbb{C}$, where $G$ is an $n\times n$ matrix with i.i.d. Gaussian entries.
Via some analytical arguments that are standard by now, it is not difficult to prove that the convergence of log potentials holds asymptotically if we remove the smallest $\epsilon n$ numbers of singular values of both $X-zI_{n}$ and $G-zI_{n}$, and thus the major task is to show the contribution of the $\epsilon n$ smallest $\sigma_{i}$’s to the sum is negligible in the limit.
For this purpose one needs to derive a lower bound for $\sigma_{min}(X-zI_{n})$. Developing a lower bound on the least singular value for $X-zI_{n}$ has been a fairly active research topic, and we refer to [MR4680362] for a review of recent breakthroughs. However, for inhomogeneous random band matrices with sub-linear bandwidth, few quantitative estimates are known. The work [jain2021circular] made the first progress in considering the block band matrix, which was followed by [tikhomirov2023pseudospectrum] who considered more general models. The bounds obtained in both papers are of the form $\sigma_{min}\geq e^{-n^{\frac{n}{d_{n}}}}$ where $d_{n}$ is the bandwidth.
Using this crude bound on $\sigma_{min}$, both [jain2021circular] and [tikhomirov2023pseudospectrum]proved circular law for certain band matrices with bandwidth $d_{n}\geq n^{\frac{32}{33}}$ (resp.$d_{n}\geq n^{\frac{33}{34}}$). To prove the circular law, the two works derived a polynomial convergence rate (see Theorem 2.3) of the empirical measure of singular values of $X_{z}X_{z}^{*}$ to $G_{z}G_{z}^{*}$ in Kolmogorov distance, and the resulting exponent $\frac{32}{33}$ arises from a careful balancing of this convergence rate and the lower bound on $\log\sigma_{min}(X_{z})$.
As one may expect, the bound $n^{\frac{32}{33}}$ appears to be far from optimal. A natural direction is to make a significant improvement of the lower bound on $\sigma_{min}(X_{z})$, as mentioned in the end of [jain2021circular], Section 1.1. However we are not able to do so at present, and we feel such improvement may not be possible for general models.
Instead, we propose that the key improvement should be that a polynomial lower bound should be available for $\sigma_{i}(X_{z})$ already for $i\sim n-n^{\tau}$ and for some specified $\tau>0$. By using this polynomial bound for $\sigma_{i}(X_{Z})$ and $i\leq n-n^{\tau}$ we can vastly improve the exponent $\frac{32}{33}$.
How do we prove this polynomial lower bound? We will show convergence of the Stieltjes transform of a dilated version (3.1) of $X_{z}$ to the deterministic limit, and the convergence holds for some mesoscopic $\Im\eta\sim n^{-\tau}$ and some $\tau>0$. Thus the Green function of (3.1) is bounded for such $\eta$, and this immediately implies a rigidity estimate on singular values of $X_{z}$, and yields a polynomial lower bound for all but a few smallest singular values of $X_{z}$.
How far can this estimate improve the bandwidth? Suppose the Green function is bounded up to the optimal scale $\Im\eta\sim(b_{n})^{-1+\epsilon}$, then we have a polynomial lower bound for all but the smallest $n^{\epsilon}\frac{n}{b_{n}}$ singular values. For these smallest ones we use the bound $|\log\sigma_{i}(X_{z})|\lesssim\frac{n}{b_{n}}\log n$ derived in Theorem 2.1 or 2.2. Then if we have $b_{n}\geq n^{\tau}$ and $\tau>\frac{1}{2}$, and taking $\epsilon$ small, we should have that the contribution of these parts vanish in the limit. This partly explains why $\tau=\frac{1}{2}$ is a critical threshold (also the threshold for Anderson localization transition), and why we cannot reach $\tau<\frac{1}{2}$ via our current method.
Unfortunately we cannot show the Green’s function of (3.1) is bounded up to $\Im\eta\gg(b_{n})^{-1}$. We can merely prove it for $\Im\eta\gg(b_{n})^{-1/5}$. At this point we mention that for a symmetric (or Hermitian) random band matrix with bandwidth $b_{n}$, local semicircle law has been established in [MR3068390] up to $\Im\eta\gg(b_{n})^{-1}$, but in our case the form of Dyson equation is different. For non-Hermitian matrix, one may also try to adapt the proof of local inhomogeneous circular law in [MR3770875] to our setting, but a crucial property of stability of Dyson equation in [MR3770875] is missing in our context. Indeed in [MR3770875] the variance profile is nowhere vanishing and has a nonzero spectral gap, but in our case the variance profile of our band matrix does not have a non-vanishing spectral gap in the $n\to\infty$ limit. Therefore we cannot use the inductive strategy in [MR3770875] to improve the scale of $\Im\eta$ and we stop at the first iteration. We also remark that in a different context, Wegner estimates have been proven in [MR3915294], [MR2525652] for certain Hermitian random band matrices which imply a polynomial lower bound on the least singular value. The proofs are however not adaptable here.
The method we take to bound the Green function of (3.1) up to $\Im\eta\gg(b_{n})^{-1/5}$ is to use a general machinery developed in [bandeira2023matrix] that yields quantitative convergence estimates of the resolvent to its free probability limit for some mesoscopic $\Im\eta$. For the non-Gaussian case we use instead the machinery in [brailovskaya2022universality]. The key point here is that [bandeira2023matrix] already yields convergence of Green function at some mesoscopic $\Im\eta$ without using stability of the Dyson equation, but does not allow to bootstrap the convergence to the optimal scale; whereas the local law approach [MR3770875] first establishes convergence for $\Im\eta\sim 1$ and then bootstraps to the optimal scale using stability of the Dyson equation (which is not avaliable here). It is very interesting to investigate whether $(b_{n})^{-1/5}$ can be improved, and whether $\eta\gg(b_{n})^{-1}$ is reachable via some recursive applications of this procedure.
Another benefit of applying the free probability approach in [bandeira2023matrix], [brailovskaya2022universality] is that we can prove circular law for a general inhomogeneous model (Definition 1.9) with doubly stochastic variance profile, whereas previous works [jain2021circular] [tikhomirov2023pseudospectrum]consider matrices that have a certain block structure on the diagonal. This is the content of Theorem 1.11 where we combine Theorem 2.2 with new quantitative convergence rates of empirical measure proved in Theorem 5.1.
Finally, we mention that in the localized regime $b_{n}\ll n^{\frac{1}{2}}$ convergence to a limiting density can still possibly be proven: there is one early work [MR2191234] which considers non-Hermitian tridiagonal matrices (the only work known so far in the regime $b_{n}\ll n^{\frac{1}{2}}$). The work essentially assumes that the off-diagonal entries have a bounded density, which enables to show convergence of the logarithm potential. The limiting density in this case is however quite complicated [MR2207197] and far from the circular law. In general, we believe that vastly different techniques are needed in the regime $b_{n}\ll n^{\frac{1}{2}}$.
2. A list of used results
In this section we outline a list of results on smallest singular values of random band matrices, and convergence rate of Stieltjes transform. All these results have been proven elsewhere.
2.1. The smallest singular value lower bound
We will need two sets of results on the smallest singular value for (non-Hermitian) random band matrix. The first set covers the case where the matrix has a block structure and in each block we have a square matrix with i.i.d. entries. For these matrices we do not require entry distribution to have a density (so that we can consider Bernoulli distribution supported on only two points).
Theorem 2.1.
[[jain2021circular], Theorem 2.1] Fix $K^{\prime}>0$. Let $X$ be a random block band matrix as in definition 1.1 with atom distribution $\xi$. Assume that $\xi$ has $p$-th moment finite for all $p\in\mathbb{N}_{+}$. Then given any $z\in\mathbb{C}$ with $|z|\leq K^{\prime}$, we have
$\mathbb{P}(s_{min}(X-zI_{n})\leq(3b_{n})^{-25\frac{n}{b_{n}}})\leq\frac{C_{\xi}}{\sqrt{3b_{n}}}$
for constant $C_{\xi}>0$ depending only on $K^{\prime}$ and the moments of $\xi$.
This theorem was proven in [jain2021circular], Theorem 2.1, under weaker moment assumption on $\xi$.
The second set of results, which we quote from Tikhomirov [tikhomirov2023pseudospectrum], is a lot more general: it does not require the matrix to be in a block form, and the location of nonzero entries can be fairly arbitrary. However it does require the entries to have a bounded density.
Theorem 2.2.
[[tikhomirov2023pseudospectrum], Theorem 3.4] Let $K,\rho_{0}>0$ be fixed. Consider a random matrix $A=V\odot W$ where $W$ is a real or complex square random matrix of size $n$, $V$ is a deterministic matrix with non-negative entries, and $\odot$ is the Hadamard (i.e., entrywise) product. Denote by
$\sigma^{*}:=\max_{i,j\leq n}|V_{i,j}|,\quad\sigma:=\max\left(\max_{j\leq n}\sqrt{\sum_{i=1}^{n}V_{i,j}^{2}},\max_{i\leq n}\sqrt{\sum_{j=1}^{n}V_{i,j}^{2}}\right).$
We further assume that the following two conditions both hold:
  1. (1)
    The entries of $W$ are independent, $K$-sub-Gaussian, having mean zero and a unit second moment.
  2. (2)
    The entries of $W$ are either real with a density bounded from above by $\rho_{0},$ or entries have independent real and imaginary parts and both have distribution densities bounded from above by $\rho_{0}$.
Then for any given $R>1$ and $\kappa\in(0,1]$, for any given $z\in\mathbb{C}$ that satisfies
$|z|>\max(\sigma^{*}n^{2\kappa},\frac{\sigma}{R}),$
we have that whenever $n$ is sufficiently large (depending on $K,R,\rho_{0},\kappa$), with probability at least $1-2n^{-2}$,
$s_{min}(A-zI_{n})\geq|z|\exp\left(-R^{2}n^{3\kappa}(\frac{\sqrt{n}\sigma^{*}}{\sigma})^{2}\right).$
It is well known (see references in [MR4680362]) that for a square matrix with i.i.d. entries, its minimal singular value is with high probability at least $n^{-C}$ for some $C>0$. However, all the bounds presented here for band matrix are super-polynomially small in $n$, and it is far from obvious whether they are sharp or not (say, is Theorem 2.2 sharp for random matrices with doubly stochastic variance profile, as mentioned in [tikhomirov2023pseudospectrum]?). In this work we do not make improvements towards these bounds, but show that they are enough for the proof of circular law up to $\gamma>\frac{5}{6}$ and illustrate that we can conceptually guess out the threshold value $\gamma=\frac{1}{2}$. From this we believe that the bounds in Theorem 2.2 are somewhat optimal for general random band matrices, and significant improvements may require very different techniques and assumptions on the matrix model.
2.2. Convergence rate of Stieltjes transform
We need two further results on a (polynomial in $n$) rate of convergence of the Stieltjes transform of random band matrix to that of the Gaussian ensemble. This is essentially derived in [jain2021circular], Section 4.2.
Consider the following empirical spectral distribution of $X_{z}:=X-zI$ and $G_{z}:=G-zI$, where $G$ is an $n\times n$ matrix with independent normalized real Gaussian distribution:
$\nu_{X_{z}}(\cdot)=\frac{1}{n}\sum_{i=1}^{n}\delta_{\sigma_{i}^{2}(X_{z})}(\cdot),$
$\nu_{G_{z}}(\cdot)=\frac{1}{n}\sum_{i=1}^{n}\delta_{\sigma_{i}^{2}(G_{z})}(\cdot).$
For two probability measures $\mu,\nu$ supported on $[0,\infty)$ we define their Kolmogorov distance via the following:
$\|\mu-\nu\|_{[0,\infty)}:=\sup_{x\geq 0}\left|\mu([0,,x])-\nu([0,,x])\right|.$ (2.1)
Then we have polynomial convergence rate of $\nu_{X_{z}}$ to $\nu_{G_{z}}$ in this distance:
Theorem 2.3.
Assume that either
  1. (1)
    $X$ is sampled from the block band matrix ensemble in Definition 1.1 and $b_{n}\geq n^{\frac{1}{2}+\epsilon}$ for some $\epsilon>0$, or
  2. (2)
    $X$ is sampled from the periodic band matrix ensemble in Definition 1.2 and $d_{n}\geq n^{\frac{1}{2}+\epsilon}$ for some $\epsilon>0$
and the atom distribution $\xi$ has finite $p$-th moment for all $p\in\mathbb{N}_{+}$. Then for any $z\in\mathbb{C}$ we can find some $\zeta>0$ (depending on $\epsilon,K^{\prime}$ and the moments of $\xi$) such that with probability $1-o(1)$, we have
$\|\nu_{X_{z}}(\cdot)-\nu_{G_{z}}(\cdot)\|_{[0,\infty)}=O(n^{-\zeta}).$ (2.2)
This theorem follows from [jain2021circular], Section 4.2 via combining equations (39), (40) therein, and the proof essentially only requires $\xi$ to have $4+\epsilon$ moments. However we need $\xi$ to have all $p$-th moments later on so we impose this restriction. Also, it is possible to derive convergence rates without assuming $b_{n}\geq n^{\frac{1}{2}+\epsilon}$: indeed we should be able co take any $b_{n}\geq n^{\epsilon}$, see [jana2017distribution] where convergence of Stieltjes transform was proven in this regime without a specified rate.
In Theorem 5.1 we will prove convergence under this much weaker condition $b_{n}\geq n^{\epsilon}$, but currently we contend ourselves with this theorem as we cannot prove circular law for $b_{n}<n^{\frac{1}{2}}$. Finally, the exponent $\zeta$ on the right hand side of (2.2) can be specified in [jain2021circular] (and the convergence rate in [jain2021circular], Lemma 4.4 can possibly be improved), but unlike [jain2021circular], in this paper the value of the exponent $\zeta$ is not particularly important to us so we do not specify $\xi$, thanks to the lower bound on smallish singular values derived in Section 3.
2.3. A linearized model
We also need a result similar to Theorem 2.1 for linearization of product matrices. This will be used to prove the product circular law for a growing number of products.
Theorem 2.4.
Let $X_{1},\cdots,X_{m_{n}}$ be independent $n\times n$ random matrices with i.i.d. entries having distribution $\xi$. Let $m_{n}$ be $n$-dependent. Assume that $\mathbb{E}[\xi]=0,\mathbb{E}[|\xi|^{2}]=1$ and $\xi$ has finite $p$-th moment for all $p\in\mathbb{N}_{+}$.
Consider the following $(nm_{n})\times(nm_{n})$ matrix
$\mathcal{L}=\begin{pmatrix}&X_{2}&&&\\ &&X_{3}&&\\ &&&\ddots&\\ &&&&X_{m}\\ X_{1}&&&&\end{pmatrix}.$ (2.3)
Then for any fixed $K^{\prime}>0$,
  1. (1)
    For any $z\neq 0$ with $|z|\leq K^{\prime}$ we have
    $\mathbb{P}(s_{min}(\mathcal{L}-zI_{n})\leq n^{-25m_{n}})\leq\frac{C_{\xi}}{\sqrt{n}}$
    for constant $C_{\xi}>0$ depending only on $K^{\prime}$ and the moments of $\xi$.
  2. (2)
    For any $z\in\mathbb{C}:|z|\leq K^{\prime}$, denote by $\mathcal{L}_{z}:=\mathcal{L}-zI_{nm_{n}}$. Assume that $m_{n}\leq n^{1-\epsilon}$ for some $\epsilon>0$, then with probability $1-o(1)$,
    $\|\nu_{\mathcal{L}_{z}}(\cdot)-\nu_{G_{z}}(\cdot)\|_{[0,\infty)}=O(n^{\zeta})$
    for some $\zeta>0$ depending only on $K^{\prime}$, $\epsilon$ and the moments of $\xi$.
While the exact statement of this theorem may be new, it can be proved by following almost verbatim the proof of Theorem 2.1 and 2.3, so we omit the details. When $m_{n}$ is independent of $n$, the first statement on the least singular value was proven in a stronger form in [MR4303893] (under finite fourth moment condition) and [MR2861673] (for subGaussian entries) with a much better quantitative estimate. See also [MR2861673] for an earlier lower bound which was polynomial in $n$ for finite $m_{n}$.
3. Stieltjes transform at intermediate scale
Throughout the section we assume that $X$ is the general inhomogeneous matrix model in Definition 1.9.
We will use the method of free probability and concentration inequalities from [bandeira2023matrix] to derive estimates on Stieltjes transform. We first work with Gaussian entries. For any $n\times n$ matrix $A\in\mathcal{M}_{n}(\mathbb{C})$ with Gaussian entries, $A$ can be written in the form
$A:=A_{0}+\sum_{i=1}^{m}A_{i}g_{i}$
where $g_{1},\cdots,g_{m}$ are i.i.d. standard real Gaussian variables of mean 0, variance one and $A_{0},A_{1},\cdots,A_{m}$ are fixed matrices in $\mathcal{M}_{n}(\mathbb{C})$. To this model $A$ we can introduce a model in free probability
$A_{\text{free}}:=A_{0}\otimes\mathbf{1}+\sum_{i=1}^{m}A_{i}s_{i},$
where $s_{1},\cdots,s_{m}$ are a free semicircular family in some $C^{*}$ algebra $\mathcal{A}$ (see Section 4.1 of [bandeira2023matrix] for more details).
We consider the dilation matrix of $X_{z}$. Define for any $z\in\mathbb{C}$ the dilation matrix
$\mathcal{Y}_{z}:=\begin{pmatrix}0&X-zI_{n}\\ X^{*}-\bar{z}I_{n}&0\end{pmatrix}.$ (3.1)
The Stieltjes transform of $\mathcal{Y}_{z}$ is defined as: for any $\eta\in\mathbb{C}_{+}:=\{z\in\mathbb{C}:\Im z>0\}$,
$\mathcal{G}_{z}(\eta):=(\mathcal{Y}_{z}-\eta I_{2n})^{-1},\quad m_{z}(\eta):=\operatorname{tr}\mathcal{G}_{z}(\eta),$
where $\operatorname{tr}$ denotes the normalized trace: for a $n\times n$ matrix $M$, $\operatorname{tr}M=\frac{1}{n}\operatorname{Tr}M$.
Our interest in the matrix $\mathcal{Y}_{z}$ stems from the fact that, a rigidity estimate on the eigenvalues of $\mathcal{Y}_{z}$ implies a rigidity estimate on singular values of $X_{z}$. We will prove a rigidity estimate for $\mathcal{Y}_{z}$ showing that not too many eigenvalues of $\mathcal{Y}_{z}$ are contained in a neighborhood of zero, which implies that except a certain amount of singular values, all other singular values of $X_{z}$ have a polynomial lower bound.
We also introduce the following free probability analogues of $\mathcal{Y}_{z}$ and $m_{z}(\eta)$. Define
$\mathcal{Y}_{z,\text{free}}:=\begin{pmatrix}0&X_{\text{free}}-zI_{n}\\ X_{\text{free}}^{*}-\bar{z}I_{n}&0\end{pmatrix},$
$\mathcal{G}_{z,\text{free}}(\eta):=(\mathcal{Y}_{z,\text{free}}-\eta I_{2n})^{-1},\quad m_{z,\text{free}}(\eta):=\operatorname{tr}\mathcal{G}_{z,\text{free}}(\eta).$
Since the matrix $X$ has doubly stochastic variance profile, we show that $m_{z,\text{free}}(\eta)$ has a deterministic expression and is independent of the underlying graph chosen. These computations have already been done in [han2024outliers], Section 3.1 and we recall them here. By [haagerup2005new], equation 1.5 and [han2024outliers], equation 3.1, the free probability Stieltjes transform $\mathcal{G}_{z,\text{free}}(\eta)$ solves the following self consistency equation
$\mathbb{E}[\mathcal{Y}_{0}\mathcal{G}_{z,\text{free}}(\eta)\mathcal{Y}_{0}]+G_{z,\text{free}}(\eta)^{-1}+\begin{pmatrix}\eta I_{N}\quad zI_{N}\\ \bar{z}I_{N}\quad\eta I_{N}\end{pmatrix}=0.$ (3.2)
This self consistency equation uniquely characterizes $\mathcal{G}_{z,\text{free}}(\eta)$ as by [helton2007operator], Theorem 2.1, for any $\eta\in\mathbb{C}_{+}$ there is a unique solution $\mathcal{G}_{z,\text{free}}(\eta)$ to this equation such that $\mathcal{G}_{z,\text{free}}(\eta)$ has a positive imaginary part. It can be checked as in [han2024outliers] that the unique solution to (3.2) with positive imaginary part is given by
$\mathcal{G}_{z,\text{free}}(\eta)=\begin{pmatrix}a(z,\eta)I_{N}&b(z,\eta)I_{N}\\ \bar{b}(z,\eta)I_{N}&c(z,\eta)I_{N}\end{pmatrix},\quad m_{z,\text{free}}(\eta)=\frac{1}{2}(a(z,\eta)+c(z,\eta)),$ (3.3)
where $a(z,\eta)$, $b(z,\eta)$ and $c(z,\eta)$ are scalar functions depending on $z$ and $\eta$. Moreover, $a$, $b$ and $c$ solve the scalar equations (see [han2024outliers], equation (3.4))
$\begin{pmatrix}c&0\\ 0&a\end{pmatrix}+\frac{1}{ac-|b|^{2}}\begin{pmatrix}c&-b\\ -\bar{b}&a\end{pmatrix}+\begin{pmatrix}\eta&z\\ \bar{z}&\eta\end{pmatrix}=0.$ (3.4)
From this computation we see that $\mathcal{G}_{z,\text{free}}(\eta)$, and hence $m_{z,\text{free}}(\eta)$, are independent of the specific variance profile for the inhomogeneous non-Hermitian matrix $X$, and in particular the Stieltjes transform $m_{z,\text{free}}(\eta)$ for $X_{z}$ coincides with the Stieltjes transform of $G_{z}$, a shifted square $n\times n$ matrix with $i.i.d.$ Gaussian entries.
The next task is to correctly identify the rate of convergence of the Stieltjes transform $m_{z}(\eta)$ towards the deterministic limit $m_{z,\text{free}}(\eta)$.
We first assume that $X$ has independent Gaussian entries, then the following theorem directly follows from applying [bandeira2023matrix], Theorem 2.8 and Corollary 4.14.
Proposition 3.1.
Assume that $X=(x_{ij})$ is a $n\times n$ random matrix. We assume that $X$ has independent Gaussian entries and a doubly stochastic variance profile $\mathbb{E}[XX^{*}]=\mathbb{E}[X^{*}X]=\mathbf{1}$. Let
$b_{n}^{-1}:=\sup_{(i,j)\in[n]^{2}}\mathbb{E}[|x_{ij}|^{2}].$
Then with probability at least $1-n^{-100}$, we have
$\|\mathcal{G}_{z}(\eta)-\mathcal{G}_{z,\text{free}}(\eta)\|_{op}\leq\frac{C}{b_{n}|\Im\eta|^{5}}+\frac{C(\log n)^{3}}{\sqrt{b}_{n}|\Im\eta|^{2}}$ (3.5)
for some universal constant $C>0$.
In particular, for any $\mathbf{c}>0$ and for any $\Im\eta\geq(b_{n})^{-1/5}n^{\mathbf{c}}$, we have with probability at least $1-n^{-100}$,
$\|\mathcal{G}_{z}(\eta)\|_{op}\leq C_{1}$ (3.6)
where $C_{1}$ is another universal constant that can be chosen uniform for all $|z|\leq 3$ and all $|\eta|\leq 10$ (we will only consider $z$ and $C$ in this range).
Proof.
The claim (3.5) directly follows from [bandeira2023matrix], Theorem 2.8 and Corollary 4.14. The bound (3.6) is then a direct consequence of the uniform upper bound on $\mathcal{G}_{z,\text{free}}$ that can be read off from the solution of the cubic equation (3.4), see for example [girko2012theory], [MR3770875]. ∎
The upper bound on the Green function (3.6) immediately implies eigenvalue rigidity at the scale $\Im\eta\gg(b_{n})^{-1/5}$.
Corollary 3.2.
Let $X$ satisfy the assumptions in Proposition 3.1. Then for any $z\in\mathbb{C}$ and any $\mathbf{c}>0$, we can find universal constants $C_{2}$ such that with probability $1-n^{-10}$, for any interval $I\subset[-5,5]$ with length
$|I|\geq(b_{n})^{-1/5}n^{\mathbf{c}},$
we have
$\{\lambda\in I:\text{$\lambda$ is an eigenvalue of $\mathcal{Y}_{z}$}\}\leq C_{2}n|I|.$
Proof.
The proof follows from a standard argument via Helffer-Sjostrand formula, and a proof for Wigner matrices can be found in [benaych2016lectures], Section 8. We only need to make two adaptations: first, we replace Stieltjes transform of semicircle law to Stieltjes transform $m_{z,\text{free}}(\eta)$; and second, we do not take $\Im\eta\sim n^{-1+\mathbf{c}}$ but only to $\Im\eta\sim(b_{n})^{-1/5}n^{\mathbf{c}}$. Finally the constant $C_{2}$ depends on a universal upper bound on $m_{z,\text{free}}(\eta)$ over $|z|\leq 3,|\eta|\leq 10$. The details are omitted ∎
The Green function estimate (3.6) also implies eigenvector delocalization via a standard argument.
Proof of Theorem 1.10, Gaussian case.
Via a continuity argument (finding grid points and using Lipschitz continuity of $\mathcal{G}_{z}$ in $z$ and $\eta$) we can show that with probability at least $1-n^{-10}$, we have the bound (3.6) holds uniformly over all $|z|\leq 3$ and $|\eta|\leq 5$.
Consider any unit eigenvector $\psi\in\mathbb{C}^{n}$ of $X$ with eigenvalue $\lambda_{\psi}$: $(X-\lambda_{\psi})\psi=0$. Then $(0,\psi)\in\mathbb{C}^{2n}$ is an eigenvector of $Y_{\lambda_{\psi}}$ with with eigenvalue 00. By classical resolvent expansion (see for example [benaych2016lectures], Theorem 2.10), for an $n\times n$ symmetric matrix $M$ with unit eigenvectors $u_{1},\cdots,u_{n}$ associated respectively to eigenvalues $\lambda_{1},\cdots,\lambda_{n}$, denoting by $G(\eta):=(M-\eta)^{-1}$ the Green function, we have for each $i,k=1,\cdots,n$,
$\Im G_{kk}(z)=\sum_{j=1}^{n}\frac{\Im\eta}{(\lambda_{j}-\lambda_{i})^{2}+(\Im\eta)^{2}}|u_{j}(k)|^{2}\geq\frac{1}{\Im\eta}|u_{i}(k)|^{2}$
Taking $\Im\eta=(b_{n})^{-1/5}n^{\mathbf{2c}}$ for some sufficiently small $\mathbf{c}>0$, this and (3.6) shows that $\|(0,\psi)\|_{\infty}\leq(b_{n})^{-1/10}n^{\mathbf{c}}$ for each eigenvector $\psi$. Thus we have $\|\psi\|_{\infty}\leq(b_{n})^{-1/10}n^{\mathbf{c}}$.
Finally, it is easy to show (see for example [bandeira2023matrix],Theorem 2.1) that with probability $1-n^{-100}$ we have $\|X\|\leq 2.5$ (where for a matrix $X$ we use $\|X\|$ to denote its operator norm), so all eigenvalues of $X$ have modulus bounded by 3.
Now we assume $X$ has a general entry distribution. In this case we will use convergence of Stieltjes transform proved in a fairly general setting in [brailovskaya2022universality].
Proposition 3.3.
Assume that $X=(x_{ij})$ is a $n\times n$ random matrix with independent entries, having mean zero and a doubly stochastic variance profile $\mathbb{E}[XX^{*}]=\mathbb{E}[X^{*}X]=\mathbf{1}$. For each $(i,j)\in[n]^{2}$ $x_{ij}$ is distributed as $b_{ij}\xi$ where $\xi$ has mean zero, variance one, finite $p$-th moment for all $p\in\mathbb{N}_{+}$ and satisfies (1.2). Set
$b_{n}^{-1}=\sup_{(i,j)\in[n]^{2}}b_{ij}^{2}$
and assume $\log_{n}b_{n}>\epsilon>0$ for some $\epsilon$. Assume that for some (small) $\mathbf{c}^{\prime}>0$,
$|\Im\eta|^{8}b_{n}\geq n^{\mathbf{c}^{\prime}}.$
Then with probability at least $1-n^{-10}$, we have
$\|\mathcal{G}_{z}(\eta)\|\leq C_{1}$ (3.7)
for a universal constant $C_{1}>0$ that can be chosen uniformly for $|z|\leq 3$ and $|\eta|\leq 5$.
Proof.
We take a truncation argument. Let $\hat{\xi}:=\xi 1_{|\xi|\leq\sqrt{b_{n}}n^{-\mathbf{c}}}$ for some sufficiently small $\mathbf{c}>0$. Let $\hat{X}$ be the random matrix constructed just as $X$ but with atom distribution $\xi$ changed to $\hat{\xi}$. Then from the moment assumption on $\xi$, we check that with probability at least $1-n^{-5}$ we have $X=\hat{X}$ and $\mathcal{G}_{z}(\eta)=\widehat{\mathcal{G}}_{z}(\eta),$ the latter being the Stieltjes transform of the dilation (3.1) of $\widehat{X}$. Let $\widehat{\mathcal{G}}_{z,\text{free}}(\eta)$ denote the Stieltjes transform defined analogously as ${\mathcal{G}}_{z,\text{free}}(\eta)$ where we replace $X_{\text{free}}$ by $\hat{X}_{\text{free}}$: $\widehat{\mathcal{G}}_{z,\text{free}}(\eta)$ can be computed similarly via (3.2) and the only difference here is that the entries of $\hat{X}$ has mean zero and a slightly smaller variance $\mathbb{E}[|\hat{\xi}|^{2}]=1-o(1)$.
Then we first apply [brailovskaya2022universality], Theorem 2.10 to the difference $\mathbb{E}[\widehat{\mathcal{G}}_{z}(\eta)-\widehat{\mathcal{G}}_{z,\text{free}}(\eta)]$ and then apply the resolvent concentration inequality in [brailovskaya2022universality], Proposition 5.6 to bound the fluctuation of $\widehat{\mathcal{G}}_{z}(\eta)$ around its mean. Analyzing the error terms implies we should take $|\Im\eta|^{8}b_{n}\geq n^{\mathbf{c}^{\prime}}$ for some $\mathbf{c}^{\prime}>0.$ Finally we take the uniform upper bound of $\widehat{G}_{z,\text{free}}(\eta)$ for $|z|\leq 3$ and $|\eta|\leq 5$.
We have been making the assumption (1.2) so that $\hat{\xi}$ is centered: this makes solving the matrix Dyson equation (3.2) a lot easier. This is the only place throughout the paper where the assumption (1.2) is used. ∎
Similarly to the Gaussian case, we have an immediate corollary with exactly the same proof that will be omitted:
Corollary 3.4.
Let $X$ satisfy the assumptions in Proposition 3.3. Then for any $z\in\mathbb{C}$ and any $\mathbf{c}>0$, we can find universal constants $C_{2}$ such that with probability $1-n^{-10}$, for any interval $I\subset[-5,5]$ with length
$|I|\geq(b_{n})^{-1/8}n^{\mathbf{c}},$
we have
$\{\lambda\in I:\text{$\lambda$ is an eigenvalue of $\mathcal{Y}_{z}$}\}\leq C_{2}n|I|.$
This also implies eigenvector delocalization bounds:
Proof of Theorem 1.10, non-Gaussian case.
The proof is the same as the Gaussian case, using Proposition 3.3 instead of Proposition 3.1. ∎
4. The circular law proof I
In this section we collect the aforementioned results on least singular value and Stieltjes transform convergence rate, to deduce convergence of the log potential and justify the circular law. This is the technique used by [jain2021circular], and earlier versions of such argument can be found in [silverstein1995empirical], [MR2567175] and [tao2008random].
To complete the proof of circular law, we shall use the replacement principle by Tao and Vu [WOS:000281425000010].
Theorem 4.1.
Assume that for each $n$, $X$ and $G$ are two ensembles of size $n\times n$ random matrices satisfying
  1. (1)
    The quantity
    $\frac{1}{n}\|G\|_{\text{HS}}^{2}+\frac{1}{n}\|X\|_{\text{HS}}^{2}$
    is bounded in probability (resp, almost surely), where $\|\cdot\|_{\text{HS}}$ denotes the Hilbert-Schmidt norm of a square matrix.
  2. (2)
    For a.e. $z\in\mathbb{C}$ with respect to Lebesgue measure, we have
    $\frac{1}{n}\log|\det(X-zI_{n})|-\frac{1}{n}\log|\det(G-zI_{n})|$
    converges to 0 in probability (resp, almost surely).
Then $\mu_{X}-\mu_{G}$ converges in probability (resp. almost surely) to zero.
Also recall that for an $n\times n$ matrix $M$, denote by $\sigma_{1}(M)\geq\sigma_{2}(M)\cdots\geq\sigma_{n}(M)$ its singular values, then
$|\det M|=\prod_{i=1}^{n}\sigma_{i}(M).$
We will also need a useful technical lemma from [jain2021circular]:
Lemma 4.2.
Consider two probability measures $\mu,\nu$ on $\mathbb{R}$ and $0<a<b$. Then
[…]
in which we define
$\|\mu-\nu\|_{[a,b]}:=\sup_{x\in[a,b]}|\mu(a,x)-\nu(a,x)|.$
Now we complete the proof of circular law in Theorem 1.5, 1.6 and 1.7.
Proofof Theorem 1.5.
First consider the Gaussian case. Take $\mathbf{c}>0$ to be a sufficiently small constant to be fixed later. We truncate the sum
$\frac{1}{n}\sum_{i=1}^{n}\log\sigma_{i}(X_{z})=\frac{1}{n}\sum_{i=1}^{n}\log\sigma_{i}(X_{z})1_{\{\sigma_{i}(X_{z})\geq(b_{n})^{-1/5}n^{\mathbf{c}}\}}+\frac{1}{n}\sum_{i=1}^{n}\log\sigma_{i}(X_{z})1_{\{\sigma_{i}(X_{z})\leq(b_{n})^{-1/5}n^{\mathbf{c}}\}}.$
By Corollary 3.2, with probability $1-n^{-100}$ there are at most $C_{2}n(b_{n})^{-1/5}n^{\mathbf{c}}$ terms in the second summation.
Then by Theorem 2.1, with probability $1-o(1)$ we have $|\log\sigma_{i}(X_{z})|\leq\frac{n}{b_{n}}\log b_{n}$ for each $i$. (We take the fact that $\|X\|\leq 3$ with high probability so that $\sigma_{1}(X_{z})$ is bounded). Thus the second sum vanishes in the limit by our assumption $\log b_{n}\geq\frac{5}{6}+\epsilon$ and upon taking $\mathbf{c}>0$ small:
$\left|\frac{1}{n}\sum_{i=1}^{n}\log\sigma_{i}(X_{z})1_{\{\sigma_{i}(X_{z})\leq(b_{n})^{-1/5}n^{\mathbf{c}}\}}\right|\leq C_{2}\frac{1}{n}n^{\mathbf{c}}\frac{n}{(b_{n})^{1/5}}\frac{n}{b_{n}}\log b_{n}=o(1),\quad n\to\infty.$ (4.1)
The same estimate in (4.1) is true if we replace $X_{z}$ by $G_{z}$ in the sum, with a much simpler proof.
Now we consider the first sum difference
$\left|\frac{1}{n}\sum_{i=1}^{n}\log\sigma_{i}(X_{z})1_{\{\sigma_{i}(X_{z})\geq(b_{n})^{-1/5}n^{\mathbf{c}}\}}-\frac{1}{n}\sum_{i=1}^{n}\log\sigma_{i}(G_{z})1_{\{\sigma_{i}(G_{z})\geq(b_{n})^{-1/5}n^{\mathbf{c}}\}}\right|$ (4.2)
$\leq 2(K_{1}+K_{2}\log n)\|\nu_{X_{z}}-\nu_{G_{z}}\|_{[0,\infty)}=o(1),\quad n\to\infty,$
where we first use Lemma 4.2 then use Theorem 2.3, and $K_{1},K_{2}$ are two universal constants. We also used $\|X_{z}\|\leq 3+|z|$ with high probability. To sum up we have shown for any $z\in\mathbb{C}$, with probability $1-o(1)$ the second criterion of Theorem 4.1 holds. The first criterion is much easier to check. Since $G$ is an i.i.d. Gaussian matrix, $\mu_{G}$ converges to the circular law, so $\mu_{X}$ also converges to circular law.
For the non-Gaussian step, we follow the same procedure but with Corollary 3.4 in place of Corollary 3.2. The only difference is now we bound, using Corollary 3.4:
$\left|\frac{1}{n}\sum_{i=1}^{n}\log\sigma_{i}(X_{z})1_{\{\sigma_{i}(X_{z})\leq(b_{n})^{-1/8}n^{\mathbf{c}}\}}\right|\leq C_{2}\frac{1}{n}n^{\mathbf{c}}\frac{n}{(b_{n})^{1/8}}\frac{n}{b_{n}}\log b_{n}=o(1),\quad n\to\infty,$ (4.3)
and the quantity is vanishing to zero thanks to the assumption $\log b_{n}\geq\frac{8}{9}+\epsilon.$
Proofof Theorem 1.6.
We follow exactly the same line as the proof of Theorem 1.5 but use Theorem 2.2, so that we can prove convergence of $\frac{1}{n}\sum_{i=1}^{n}\log\sigma_{i}(X_{z})$ to $\frac{1}{n}\sum_{i=1}^{n}\log\sigma_{i}(G_{z})$ for any $z\in\mathbb{C}$, $z\neq 0$ (note that the case $z=0$ is excluded by Theorem 2.2). This is sufficient to prove circular law by Theorem 4.1.
Proofof Theorem 1.7.
We first show that proving the circular law for $\mathcal{M}$ is equivalent to proving the circular law for $\mathcal{L}$ as defined in (2.3): this is a standard linear algebra exercise. Indeed, if $\lambda$ is an eigenvalue of $\mathcal{M}$, then $\lambda^{m_{n}}$ is an eigenvalue of $X_{1}\cdots X_{m_{n}}$ and vice versa. Meanwhile, by [MR4076784], Proposition 4.1, eigenvalues of $X_{1}\cdots X_{m_{n}}$ and eigenvalues of $\mathcal{L}^{m_{n}}$ are equal with multiplicity $m_{n}$.
Thus it suffices to prove circular law for $\mathcal{L}$. For this we follow exactly the proof of Theorem 1.5 in the non-Gaussian case where we use Theorem 2.4 for the model $\mathcal{L}$. We require that $n\geq(nm_{n})^{\frac{8}{9}+\epsilon}$, so that $m_{n}\leq n^{\frac{1}{8}-\epsilon^{\prime}}$ for some $\epsilon^{\prime}>0$. The details are omitted. ∎
5. The circular law proof II
In this section we outline proof of the circular law for the very general model 1.9. All the proof ingredients are available except a convergence rate of the empirical measure, which we will prove next.
Theorem 5.1.
Let $X$ be defined as in Definition 1.9 but with $b_{n}\geq n^{\epsilon}$ for some small $\epsilon>0$, and with $\xi$ having Gaussian distribution. Then for any fixed $z\in\mathbb{C}$ we can find $\zeta>0$ such that with probability $1-o(1)$, we have
$\|\nu_{X_{z}}(\cdot)-\nu_{G_{z}}(\cdot)\|_{[0,\infty)}=O(n^{-\zeta}).$ (5.1)
In this result we do not need $b_{n}\geq n^{\frac{5}{6}+\epsilon}$ but merely $b_{n}\geq n^{\epsilon}$. For some specific band matrix models the convergence has been proven in [jana2017distribution], [jain2021circular]. However, no proof was previously given for this very general case, i.e. assuming only a doubly stochastic variance profile. Also, the assumption that $\xi$ has Gaussian distribution is made just to simplify the proof, and exactly the same holds for general $\xi$ in Definition 1.9 where we use the machinery from [brailovskaya2022universality] instead of [bandeira2023matrix].
Proof of Theorem 5.1.
Let $\widetilde{G}_{z}$ denote the dilation matrix
$\widetilde{G}_{z}:=\begin{pmatrix}0&G-zI\\ G^{*}-\bar{z}I&0\end{pmatrix}.$
Observing that $\lambda$ is an eigenvalue of $X_{z}X_{z}^{*}$ (resp. $G_{z}G_{z}^{*}$) if and only if ±$\pm\sqrt{\lambda}$ are both eigenvalues of $\mathcal{Y}_{z}$ (resp. $\widetilde{G}_{z}$), we see that to prove (5.1), it suffices to prove
$\|\mu_{\mathcal{Y}_{z}}(\cdot)-\mu_{\widetilde{G}_{z}}(\cdot)\|_{(-\infty,\infty)}=O(n^{-\zeta}),$ (5.2)
where we recall that $\mu_{X}$ denotes the empirical eigenvalue distribution: it differs from $\nu_{X}$ by considering the eigenvalues instead of considering squared singular values, but this change is compatible with definition of the Kolmogorov distance up to a universal constant. In other words, a bound on $\|\mu_{\mathcal{Y}_{z}}(\cdot)-\mu_{\widetilde{G}_{z}}(\cdot)\|_{(-\infty,\infty)}$ immediately implies a bound on $\|\nu_{\mathcal{Y}_{z}}(\cdot)-\nu_{\widetilde{G}_{z}}(\cdot)\|_{[0,\infty)}$, which by definition equals $\|\nu_{X_{z}}(\cdot)-\nu_{G_{z}}(\cdot)\|_{[0,\infty)}$.
The benefit of this reduction is that both $\mathcal{Y}_{z}$ and $\widetilde{G}_{z}$ have independent entries modulo symmetry. To prove (5.2), we show the Stieltjes transform of $\mathcal{Y}_{z}$ converges to the Stieltjes transform of $\widetilde{G}_{z}$ at a polynomial rate. Recall that $m_{z}(\eta):=\operatorname{tr}\mathcal{G}_{z}(\eta)$ is the Stieltjes transform of $\mathcal{Y}_{z}$, and $m_{z,\text{free}}(\eta)$ is the Stieltjes transform of $\mathcal{Y}_{z,\text{free}}$. Here the expression $m_{z,\text{free}}(\eta)$ has been explicitly computed in (3.3), (3.4).
If we let $g_{z,\text{free}}(\eta)$ denote the Stieltjes transform of $\widetilde{G}_{z,\text{free}}$ then we have $m_{z,\text{free}}(\eta)=g_{z,\text{free}}(\eta)$ as they both satisfy the defining relations (3.3), (3.4). Hence in the following we use $\mu_{z}(\cdot)$ to denote both $\mu_{\mathcal{Y}_{z},\text{free}}(\cdot)$ and $\mu_{\widetilde{G}_{z},\text{free}}(\cdot)$. By [MR2663633], Lemma 3.1 and Remark 3.1, $\mu_{z}(\cdot)$ has a bounded support and bounded density: for each $x\in\mathbb{R},y\geq 0:$
$|\mu_{z}((-\infty,x+y))-\mu_{z}((-\infty,x))|\leq y.$ (5.3)
Using Theorem 2.8 and Corollary 4.14 of [bandeira2023matrix], we deduce that for any $z\in\mathbb{C}$ and any $\eta$ with $\Im\eta>0$, with probability $1-n^{-100}$ we have
$|m_{z}(\eta)-m_{z,\text{free}}(\eta)|\leq\frac{(\log n)^{5}}{b_{n}|\Im\eta|^{5}}.$ (5.4)
Fix an arbitrarily large $A>0$ and Denote by $\mathcal{D}_{A}$ the region
$\mathcal{D}_{A}:=\{z\in\mathbb{C}:-A\leq\Re z\leq A,(b_{n})^{-1/6}\leq\Im z\leq 1\}.$
Using the Lipschitz continuity of $m_{z}(\eta)$ and $g_{z}(\eta)$ in $\eta$, we can upgrade the convergence in (5.4) to be uniform over $\eta\in\mathcal{D}_{A}$, with probability at least $1-n^{-10}$.
In the following we take $K>0$ sufficiently large such that $\mu_{\mathcal{Y}_{z}}(\mathbb{R}\setminus[-K,K])=0$ with probability $1-o(1)$, and $\mu_{\mathcal{Y}_{z,\text{free}}}(\mathbb{R}\setminus[-K,K])=0$. Also take some $a>0$ sufficiently large. Now we use [MR2567175], Corollary B.15 to derive, for some $A>0$ large depending on $K$ and $a$, (writing $\xi=\theta+i\tau$)
$\|\mu_{\mathcal{Y}_{z}}(\cdot)-\mu_{z}(\cdot)\|_{(-\infty,\infty)}$
[…]
for some $C$ depending only on $A$, $a$ and $K$. Now we set $\tau=(b_{n})^{-1/6}$. To bound the first term on the second line we use (5.4), and to bound the second term on the second line use (5.3). Doing a same computation for $\widetilde{G}_{z}$ completes the proof of (5.2).
Now we are ready to complete the proof of Theorem 1.11.
Proofof Theorem 1.11.
We have the same estimate (4.1) for any $z\neq 0$ with probability $1-o(1)$: this follows from applying Theorem 2.2.
We also have the same estimate (4.2) with probability $1-o(1)$ which follows from applying Theorem 5.1. This completes the proof. ∎
Funding
The author is supported by a Simons Foundation Grant (601948, DJ)
[…]