### v2

# The circular law for random band matrices: improved bandwidth for general models

Abstract.
We consider the convergence of the ESD for non-Hermitian random band matrices with independent entries to the circular law, which is the uniform measure on the unit disk in the center of the complex plane. We assume that the bandwidth of the matrix scales like $n^{\gamma}$ for some $\gamma\in(0,1]$, where $n$ is the matrix size, and the variance profile of the matrix is only assumed to be doubly stochastic with no additional assumption on its specific mixing properties. We prove that the circular law limit holds either (1) when $\gamma>\frac{5}{6}$ and the entries are independent Gaussians, (2) or when $\gamma>\frac{8}{9}$ and the entries are independent subgaussian random variables. This new threshold improves the previous threshold $\gamma>\frac{32}{33}$ which was only proven for block band matrices and periodic band matrices. After the initial version of this paper, the author further extended the range of circular law for much smaller values of $\gamma$ in [han2025circular1] and [han2025circular2] when the variance profile has specific mixing properties, but not for an arbitrary doubly stochastic variance profile. Thus the main contribution of this paper is the circular law for a genuine power law bandwidth for any doubly stochastic variance profile. We also prove an extended form of product circular law with a growing number of matrices. Weak delocalization estimates on eigenvectors are also derived. The new technical input is new polynomial lower bounds on some intermediate small singular values, and this estimate does not depend on the specific structure of the variance profile beyond the fact that it is doubly stochastic.
1. Introduction
For an $n\times n$ matrix $A$ with eigenvalues $\lambda_{1}(A),\cdots,\lambda_{n}(A)\in\mathbb{C}$ we denote its empirical measure of eigenvalues by
$\mu_{A}:=\frac{1}{n}\sum_{i=1}^{n}\delta_{\lambda_{i}(A)},$
where $\delta_{z}$ is the Dirac measure at point $z$.
For a symmetric random matrix $A$, the convergence of $\mu_{A}$ to the semicircle distribution under mild assumptions dates back to the celebrated work of Wigner [wigner1958distribution]. The spectrum of symmetric random matrices can essentially be determined via the method of moments. For non-Hermitian random matrix $A$, the convergence of $\mu_{A}$ does not follow from method of moments arguments (see for example [MR2908617]), and fundamentally new ideas are needed. When all the entries of $A=(a_{ij})_{1\leq i,j\leq n}$ are i.i.d. with mean zero and variance one, the circular law theorem, i.e. the convergence of $\mu_{A}$ to the uniform distribution on the unit disk in the complex plane, is one of the fundamental theorems in modern random matrix theory. This circular law was proven under increasing generality in a series of works including [girkoarticle], [MR1428519], [tao2008random], [WOS:000281425000010], and we refer to the references in these works and [MR2908617] for a more complete list. Beyond dense matrices with i.i.d. entries, the circular law was subsequently proven for sparse i.i.d. matrices [MR3980923], [MR3945840] [sah2023limiting], for sparse regular digraphs [MR4195739], and for some random matrices with inhomogeneous variance profile [MR3878135]. A more complete list of references can be found in the review [MR4680362].
Despite all the recent efforts in proving the circular law, effective techniques for structured non-Hermitian random matrices with inhomogeneous variance profile are still out of reach. In these models, a large part of entries are always set to be zero. (Note that for dense inhomogeneous matrices [MR3857860], [MR3878135] developed a general technique to prove circular law, but the approach becomes ineffective for sparse inhomogeneous models). The guiding examples of sparse inhomogeneous matrices are random band matrices, which we will define shortly. Their Hermitian analogue, namely Hermitian Random band matrices, play a fundamental role in modern mathematical physics. They serve as prototypical models of Anderson transition: when the bandwidth scales like $n^{\gamma}$ for $\gamma>\frac{1}{2}$, the eigenvectors are expected to be delocalized and eigenvalues have GOE local statistics; whereas a localized phase and Poisson statistics should occur when $\gamma<\frac{1}{2}$. Some very recent progress on this topic can be found in [yau2025delocalization] , [erdHos2025zigzag], [dubova2025delocalization], [dubova2025delocalization3d], [drogin2025localization],[shcherbina2025characteristic] and the references therein.
In this work we study non-Hermitian versions of random band matrix, where even global universality, i.e. the convergence of the empirical spectral density to the circular law, remains largely unproven. Two recent papers [jain2021circular] [tikhomirov2023pseudospectrum] justified the circular law for certain non-Hermitian band matrices with bandwidth $n^{\gamma}$ with $\gamma$ sufficiently close to 1: more precisely they require $\gamma>\frac{32}{33}$. The main contribution of this paper is to consider narrower bandwidth and expand the variety of models where global universality holds: we show $\gamma>\frac{5}{6}$ is enough for the circular law limit for Gaussian entries whenever the variance profile is doubly stochastic.
1.1. Circular law for general models
Definition 1.1.
A real or complex random variable $\xi$ of mean zero, variance one is said to be $K$-subGaussian if
$\mathbb{E}\exp(|\xi|^{2}/K^{2})\leq 2.$
A mean 0, variance 1 random variable $\zeta$ has finite $p$-th moment for all $p$, if for all $p\in\mathbb{N}_{+}$ we can find $C_{p}>0$ such that
$\mathbb{E}[|\xi|^{p}]\leq C_{p}<\infty,\quad\forall p\in\mathbb{N}_{+}.$
We consider the following very general definition of inhomogeneous matrices:
Definition 1.2 (General model).
We define an ensemble of inhomogeneous random matrices as follows: $X=(x_{ij})$ is a $n\times n$ random matrix with independent entries. For each $(i,j)\in[n]^{2}$, $x_{ij}$ is distributed as $b_{ij}\xi$, where $b_{ij}\geq 0$ and $\xi$ has mean zero, variance one, and is $K$-subgaussian for some fixed $K>0$. The bandwidth of this general matrix model is defined by
$b_{n}^{-1}=\sup_{(i,j)\in[n]^{2}}b_{ij}^{2}.$
We assume that $X$ has a doubly stochastic variance profile:
$\mathbb{E}[XX^{*}]=\mathbb{E}[X^{*}X]=I_{n}.$
In other words, we assume that
$\sum_{j\in[n]}b_{ij}^{2}=1\quad\forall i\in[n],\quad\sum_{i\in[n]}b_{ij}^{2}=1\quad\forall j\in[n].$
The generality of this definition is already illustrated by the following subclasses of models: consider $G_{n}=([n],\mathcal{E})$ a $b_{n}$- regular directed graph on $n$ vertices. Then we set $b_{ij}=\frac{1}{\sqrt{b_{n}}}$ if $(i,j)\in\mathcal{E}$ and $b_{ij}=0$ otherwise. Observe that the graph $G_{n}$ can be arbitrary and we never require the structure of $G_{n}$ to be sufficiently pseudo-random.
Example 1.3.
(Well-mixing variance profiles) We first review some typically used variance profiles for a random band matrix in its non Hermitian form, that have rapid mixing properties. The first example is the following defined periodic band matrix. Let $X$ be an $n\times n$ matrix. We say $X$ is sampled from a block band matrix ensemble with bandwidth $b_{n}$ if it has the form
$X=\begin{pmatrix}D_{1}&U_{2}&&&T_{m}\\ T_{1}&D_{2}&U_{3}&&\\ &T_{2}&D_{3}&\ddots&\\ &&\ddots&\ddots&U_{m}\\ U_{1}&&&T_{m-1}&D_{m}\end{pmatrix}$ (1.1)
where the unfilled sites are set zero. The blocks $D_{1},U_{1},T_{1},\cdots,D_{m},U_{m},T_{m}$ are $b_{n}/3\times b_{n}/3$ independent random matrices with i.i.d. entries having the same distribution $\frac{1}{\sqrt{b_{n}}}\xi$, where $\xi$ is a standard Gaussian.
The second example is periodic band matrices defined as follows. Let $X$ be an $n\times n$ square matrix. We say $X$ is sampled from a periodic band matrix model with bandwidth $b_{n}$ if $x_{ij}=0$ for any $\frac{b_{n}-1}{2}<|i-j|<n-\frac{b_{n}-1}{2}$, and $x_{ij}$ are i.i.d. copies of $\frac{1}{\sqrt{b_{n}}}\xi$ if $|i-j|\leq\frac{b_{n}-1}{2}$ or $|i-j|\geq n-\frac{b_{n}-1}{2}$. Here $\xi$ is a standard gaussian and we assume $\frac{b_{n}-1}{2}$ is an integer.
In some other concrete settings of applications, and as a conceptual question itself, it is of major interest to study variance profiles that do not have good mixing properties. For example, we can consider the following variance profiles that are vertical translations of the variance profile of a block or periodic band matrix:
Example 1.4.
(Variance profiles without good mixing) For a given bandwidth $b_{n}$ that is divisible by $n$, we can consider the following two variance profiles: (1) shifted periodic band matrix where for example we take $b_{ij}=\frac{1}{\sqrt{b_{n}}}\mathbf{1}(\min(|i-j-3b_{n}|,|n-i+j+3b_{n}|)\leq\frac{b_{n}-1}{2})$, this corresponds to right-shifting the variance of a periodic band matrix by $3b_{n}$; and (2) variance profile from matrix product, where $b_{ij}=\frac{1}{\sqrt{b_{n}}}\mathbf{1}(\lfloor\frac{j-1}{b_{n}}\rfloor-\lfloor\frac{i-1}{b_{n}}\rfloor=1\text{ or }1-\frac{n}{b_{n}})$, which corresponds to preserving only the $U_{i}$ blocks of $X$ in (1.1), setting all other blocks of $X$ to be zero and taking a suitable rescaling to be doubly stochastic.
The main results of this paper are as follows.
Theorem 1.5 (Circular law for general model).
Let $X$ be the general inhomogeneous random matrix model defined in Definition 1.2, and we assume that
  1. (1)
    Either $\xi$ has Gaussian distribution, and $b_{n}\geq n^{\frac{5}{6}+\epsilon}$ for some $\epsilon>0$,
  2. (2)
    or $\xi$ is $K$-subgaussian for some $K>0$, and $\xi$ is either real-valued with a distributional density on $\mathbb{R}$ bounded by some $L>0$, or $\xi$ has independent real and imaginary parts, both having a distributional density bounded by some $L>0$.
Then the ESD $\mu_{X}$ of $X$ converges in probability to the circular law.
Remark 1.6.
Two recent papers by the author, [han2025circular1] and [han2025circular2], show that when the variance profile has good mixing properties, then $\frac{5}{6}+\epsilon$ is not the optimal threshold of circular law. Thus Theorem 1.5 is suboptimal for block band matrices and periodic band matrices and the result is superseded in [han2025circular1]. However, neither of these papers cover an arbitrary variance profile: for instance, they do not cover all the examples in Example 1.4. Thus Theorem 1.5 remains the first result of circular law for these very general variance profiles, and is a strong justification that arbitrary normalized variance profiles can have the circular law limit in this sublinearly growing bandwidth regime.
Via a further application of the linearization trick, we prove a circular law for the product of a growing number of i.i.d. matrices:
Theorem 1.7.
Let $X_{1},\cdots,X_{m_{n}}$ be independent $n\times n$ random matrices with i.i.d. entries having distribution $\frac{1}{\sqrt{n}}\xi$. Let $m=m_{n}$ be $n$-dependent and such that $m_{n}\leq n^{\frac{1}{8}-\epsilon}$ for some $\epsilon>0$. Assume that $\mathbb{E}[\xi]=0,\mathbb{E}[|\xi|^{2}]=1$ and $\xi$ has finite $p$-th moment for all $p\in\mathbb{N}_{+}$. Further assume that $\xi$ is either real-valued with distributional density bounded by $L>0$, or $\xi$ has independent real and imaginary parts, at least one of them has distributional density bounded by $L>0$. Denote by
$\mathcal{M}:=(X_{1}\cdots X_{m_{n}})^{\frac{1}{m_{n}}}.$
Then in the limit $n\to\infty$, we have that the empirical measure $\mu_{\mathcal{M}}$ converges in probability to the circular law.
Note that in this result, we assume a slightly weaker moment assumption on $\xi$.
Remark 1.8.
It is not clear whether or not we can consider a discrete entry law $\xi$ in Theorem 1.7, due to instability of the linearized matrix $\mathcal{L}$ (2.2). This instability of $\mathcal{L}$ disqualifies us from generalizing the proof of [jain2021circular], Theorem 2.1 to this setting. When $m_{n}$ is fixed and independent of $n$, the circular law of $\mathcal{M}$ has been proved in [MR2861673] under a $2+\eta$-th moment condition. That is, the eigenvalues of $X_{1}\cdots X_{m_{n}}$ converges to the so-called $m_{n}$-fold product circular law. A local circular law has also been proven for product matrices [MR3622892], and subsequently universality of linear statistics for i.i.d matrix products are derived [MR4112718]. Via adapting the proof in [jain2021circular], we can possibly relax the moment condition on $\xi$ in Theorem 1.7 to, say, assuming only a finite six-th moment, at the cost of allowing a much smaller rate of growth of $m_{n}$. As the focus here is on large $m_{n}$, we sacrifice with the moment assumptions.
Optimally, one expects to prove the circular law for all $m_{n}\leq n^{1-\epsilon}$, but this is beyond the techniques in this work. Indeed, the proof of Theorem 1.7 relies on the proof of circular law for the linearization matrix (2.2) of this matrix product, but this linearization matrix does not satisfy the specific variance properties stated in [han2025circular1] because the variance profile does not have rapid mixing property (see Example 1.4), so it precludes the regime $m_{n}\sim n^{1-c}$ stated in [han2025circular1]. The threshold $m_{n}\sim n$ has reached a lot of recent attention. Informally, $m_{n}\ll n$ is the free probability regime where universality is expected to hold for the product matrix; $m_{n}\sim n$ is the critical transition regime and $m_{n}\gg n$ is the ergodic regime (and there should be no universality). We refer to [MR4401507], [MR4421171] and [MR4580535] for integrable probability perspectives on product random matrices with simultaneously growing $m_{n}$ and $n$, and to [MR4268303] for an ergodic theory perspective.
1.2. Weak delocalization estimates
A major component of our proof is to upper bound the Green function $G(z)$ of a dilated version (3.1) of $X$, at some scales $\Im z\ll 1$. We cannot reach the optimal scale $\Im z\sim n^{-1+\epsilon}$ and are indeed very far from that, but such a bound appears to be new and implies interesting delocalization estimates.
We introduce the most general model where our delocalization statement is valid.
Theorem 1.9 (Weak delocalization).
Consider a general random matrix ensemble in Definition (1.2). Assume that $b_{n}\geq n^{\epsilon}$ for some sufficiently small $\epsilon>0$. Then for any (sufficiently small) $\mathbf{c}>0$ there holds
  1. (1)
    If $\xi$ has (real or complex) Gaussian distribution. then with probability at least $1-n^{-10}$, all eigenvectors $\psi$ of $X$ with unit $L^{2}$ norm satisfy
    $\|\psi\|_{\infty}\leq(b_{n})^{-1/10}n^{\mathbf{c}}.$
  2. (2)
    For a general mean zero, variance 1 random variable $\xi$ having all $p$-th moment finite, then with probability at least $1-n^{-10}$, all eigenvectors $\psi$ of $X$ with unit $L^{2}$ norm satisfy
    $\|\psi\|_{\infty}\leq(b_{n})^{-1/16}n^{\mathbf{c}}.$
Remark 1.10.
[…]
1.3. Main ideas and outline
We follow the classical Hermitization trick of Girko to establish the circular law, which we refer to [MR2908617] for a thorough introduction and historical review.
For an $n\times n$ matrix $A$ we denote by $\sigma_{1}(A)\geq\sigma_{2}(A)\cdots\geq\sigma_{n}(A)$ the list of singular values of $A$, arranged in decreasing order.
Thanks to the replacement principle by Tao and Vu (see Theorem 4.1), the main technical step to prove circular law is to show convergence of the logarithmic sums of singular values of shifted matrix $X_{z}:=X-zI_{n}$, $\sum_{i=1}^{n}\log\sigma_{i}(X-zI_{n})$, towards $\sum_{i=1}^{n}\log\sigma_{i}(G-zI_{n})$ for a.e. $z\in\mathbb{C}$, where $G$ is an $n\times n$ matrix with i.i.d. Gaussian entries.
Via some analytical arguments that are standard by now, it is not difficult to prove that the convergence of log potentials holds asymptotically if we remove the smallest $\epsilon n$ numbers of singular values of both $X-zI_{n}$ and $G-zI_{n}$, and thus the major task is to show the contribution of the $\epsilon n$ smallest $\sigma_{i}$’s to the sum is negligible in the limit.
For this purpose one needs to derive a lower bound for $\sigma_{min}(X-zI_{n})$. Developing a lower bound on the least singular value for $X-zI_{n}$ has been a fairly active research topic, and we refer to [MR4680362] for a review of recent breakthroughs. However, for inhomogeneous random band matrices with sub-linear bandwidth, few quantitative estimates are known. The work [jain2021circular] made the first progress in considering the block band matrix, which was followed by [tikhomirov2023pseudospectrum] who considered more general models. The bounds obtained in both papers are of the form $\sigma_{min}\geq e^{-n^{\frac{n}{d_{n}}}}$ where $d_{n}$ is the bandwidth.
[…]
As one may expect, the bound $n^{\frac{32}{33}}$ appears to be far from optimal. A natural direction is to make a significant improvement of the lower bound on $\sigma_{min}(X_{z})$, as mentioned in the end of [jain2021circular], Section 1.1. However we are not able to do so at present, and we feel such improvement may not be possible for general models.
Instead, we propose that we can make an improvement by obtaining a polynomial lower bound for small-ish singular values $\sigma_{i}(X_{z})$ already for $i\sim n-n^{\tau}$ and for some specified $\tau>0$. The smaller the value of $\tau$, the better convergence for circular law would be. By using this polynomial bound for $\sigma_{i}(X_{z})$ and $i\leq n-n^{\tau}$, we can quantitatively lower the exponent $\frac{32}{33}$ to a much smaller value.
How do we prove this polynomial lower bound? We will show convergence of the Stieltjes transform of a dilated version (3.1) of $X_{z}$ to the deterministic limit, and the convergence holds for some mesoscopic $\Im\eta\sim n^{-\tau}$ and some $\tau>0$. Thus the Green function of (3.1) is bounded for such $\eta$, and this immediately implies a rigidity estimate on singular values of $X_{z}$, and yields a polynomial lower bound for all but a few smallest singular values of $X_{z}$.
How far can this estimate improve the bandwidth? Suppose the Green function is bounded up to the scale $\Im\eta\sim(b_{n})^{-1+\epsilon}$, then we can have a polynomial lower bound for all but the smallest $n^{\epsilon}\frac{n}{b_{n}}$ singular values. For these smallest ones we use the bound $|\log\sigma_{i}(X_{z})|\lesssim\frac{n^{1+\epsilon}}{b_{n}}$ for some $\epsilon>0$, which directly follows from Theorem 2.2 or 2.1. Then if we have $b_{n}\geq n^{\tau}$ and $\tau>\frac{1}{2}$, and taking $\epsilon$ small, we should have that the contribution of these parts vanish in the limit. This would lead to a threshold $\tau=\frac{1}{2}$.
By the time the paper was initially written, it was not clear how to show the Green’s function of (3.1) is bounded up to $\Im\eta\gg(b_{n})^{-1}$. For an arbitrary variance profile, we can merely prove it for $\Im\eta\gg(b_{n})^{-1/5}$. To compare, for a symmetric (or Hermitian) random band matrix with bandwidth $b_{n}$, local semicircle law can be easily established up to $\Im\eta\gg(b_{n})^{-1}$ (see [MR3068390]), but this result requires some mixing properties of the variance profile. After the initial version of this paper was completed, the author succeeded in adapting the proof of local inhomogeneous circular law in [MR3770875] to our non-Hermitian setting in [han2025circular1], under the assumption that the variance profile has certain mixing properties. This achieves the threshold $\tau=\frac{1}{2}$. However, if we consider an arbitrary doubly stochastic variance profile, we cannot use the inductive strategy in [MR3770875] to improve the scale of $\Im\eta$ and thus we do not succeed in proving any local law from this method. We also remark that in a different context, Wegner estimates have been proven in [MR3915294], [MR2525652] for certain Hermitian random band matrices which imply a polynomial lower bound on the least singular value. The proofs are however not adaptable in our model because there are no random potentials on the diagonal in the Hermitization of our non-Hermitian band matrix.
To handle an arbitrary variance profile, our method to bound the Green function of (3.1) up to $\Im\eta\gg(b_{n})^{-1/5}$ is to use a general machinery developed in [bandeira2023matrix] that yields quantitative convergence estimates of the resolvent to its free probability limit for some mesoscopic $\Im\eta$. For the non-Gaussian case we use instead the machinery in [brailovskaya2022universality]. The key point here is that [bandeira2023matrix] already yields convergence of Green function at some mesoscopic $\Im\eta$ via concentration inequalities, without using stability of the matrix Dyson equation. This is very helpful because the stability of matrix Dyson equation is not guaranteed for an arbitrary variance profile. Another benefit of applying the free probability approach in [bandeira2023matrix], [brailovskaya2022universality] is that it applies immediately to any general model (Definition 1.2) with doubly stochastic variance profile, whereas [jain2021circular] only considers matrices that have a certain block structure on the diagonal.
Although the threshold of this paper was recently surpassed by [han2025circular1] and [han2025circular2] for specific models, its applicability to any doubly stochastic variance profile remains attractive and unique. Moreover, the idea highlighted in this paper, that improved rigidity estimates strengthen the circular law limit, is also one of the prevailing themes in [han2025circular1] and [han2025circular2]: indeed, both works still use the proof idea of this paper (via free probability) in certain intermediate steps, and the reason why these papers have sharper thresholds is that they combine many new ideas and techniques into the proof.
2. A list of used results
In this section we outline a list of results on smallest singular values of random band matrices, and convergence rate of Stieltjes transform. All these results have been proven elsewhere.
2.1. The smallest singular value lower bound
We will need a result on the smallest singular value for (non-Hermitian) random band matrix by Tikhomirov [tikhomirov2023pseudospectrum].
Theorem 2.1.
[[tikhomirov2023pseudospectrum], Theorem 3.4] Let $K,\rho_{0}>0$ be fixed. Consider a random matrix $A=V\odot W$ where $W$ is a real or complex square random matrix of size $n$, $V$ is a deterministic matrix with non-negative entries, and $\odot$ is the Hadamard (i.e., entrywise) product. Denote by
$\sigma^{*}:=\max_{i,j\leq n}|V_{i,j}|,\quad\sigma:=\max\left(\max_{j\leq n}\sqrt{\sum_{i=1}^{n}V_{i,j}^{2}},\max_{i\leq n}\sqrt{\sum_{j=1}^{n}V_{i,j}^{2}}\right).$
We further assume that the following two conditions both hold:
  1. (1)
    The entries of $W$ are independent, $K$-sub-Gaussian, having mean zero and a unit second moment.
  2. (2)
    The entries of $W$ are either real with a density bounded from above by $\rho_{0},$ or entries have independent real and imaginary parts and both have distribution densities bounded from above by $\rho_{0}$.
Then for any given $R>1$ and $\kappa\in(0,1]$, for any given $z\in\mathbb{C}$ that satisfies
$|z|>\max(\sigma^{*}n^{2\kappa},\frac{\sigma}{R}),$
we have that whenever $n$ is sufficiently large (depending on $K,R,\rho_{0},\kappa$), with probability at least $1-2n^{-2}$,
$s_{min}(A-zI_{n})\geq|z|\exp\left(-R^{2}n^{3\kappa}(\frac{\sqrt{n}\sigma^{*}}{\sigma})^{2}\right).$
It is well known (see references in [MR4680362]) that for a square matrix with i.i.d. entries, its minimal singular value is with high probability at least $n^{-C}$ for some $C>0$. However, all the bounds presented here for band matrix are super-polynomially small in $n$, and it is far from obvious whether they are sharp or not (say, is Theorem 2.1 sharp for random matrices with doubly stochastic variance profile, as mentioned in [tikhomirov2023pseudospectrum]?). In this work we do not make improvements towards these bounds, but show that they are enough for the proof of circular law up to $\gamma>\frac{5}{6}$ and illustrate that we can conceptually guess out the threshold value $\gamma=\frac{1}{2}$. From this we believe that the bounds in Theorem 2.1 are somewhat optimal for general random band matrices, and significant improvements may require very different techniques and assumptions on the matrix model.
2.2. Convergence rate of Stieltjes transform
We need two further results on a (polynomial in $n$) rate of convergence of the Stieltjes transform of random band matrix to that of the Gaussian ensemble. This is essentially derived in [jain2021circular], Section 4.2.
Consider the following empirical spectral distribution of $X_{z}:=X-zI$ and $G_{z}:=G-zI$, where $G$ is an $n\times n$ matrix with independent normalized real Gaussian distribution:
$\nu_{X_{z}}(\cdot)=\frac{1}{n}\sum_{i=1}^{n}\delta_{\sigma_{i}^{2}(X_{z})}(\cdot),$
$\nu_{G_{z}}(\cdot)=\frac{1}{n}\sum_{i=1}^{n}\delta_{\sigma_{i}^{2}(G_{z})}(\cdot).$
For two probability measures $\mu,\nu$ supported on $0,\infty)$ we define their Kolmogorov distance via the following:
$\|\mu-\nu\|_{[0,\infty)}:=\sup_{x\geq 0}\left|\mu([0,,x])-\nu([0,,x])\right|.$ (2.1)
2.3. A linearized model
We also need a least singular value result similar to [jain2021circular], Theorem 2.1 for linearization of product matrices. This will be used to prove the product circular law for a growing number of products.
Theorem 2.2.
Let $X_{1},\cdots,X_{m_{n}}$ be independent $n\times n$ random matrices with i.i.d. entries having distribution $\frac{1}{\sqrt{n}}\xi$. Let $m_{n}$ be $n$-dependent. Assume that $\mathbb{E}[\xi]=0,\mathbb{E}[|\xi|^{2}]=1$ and $\xi$ has finite $p$-th moment for all $p\in\mathbb{N}_{+}$. Moreover, assume that $\xi$ is either real with distributional density bounded by $L>0$, or have independent real and complex parts and one of them has distributional density bounded by $L>0$.
Consider the following $(nm_{n})\times(nm_{n})$ matrix
$\mathcal{L}=\begin{pmatrix}&X_{2}&&&\\ &&X_{3}&&\\ &&&\ddots&\\ &&&&X_{m}\\ X_{1}&&&&\end{pmatrix}$. (2.2)
Then for any fixed $K^{\prime}>0$, for any $z\neq 0$ with $|z|\leq K^{\prime}$ we have
$\mathbb{P}(s_{min}(\mathcal{L}-zI_{nm_{n}})\leq\min(|z|,1/|z|)^{2m_{n}}n^{-25m_{n}})\leq\frac{C_{\xi}}{\sqrt{n}}$
for constant $C_{\xi}>0$ depending only on $K^{\prime}$ and the moments of $\xi$.
Theorem 2.2 is proven in Section 5.
Remark 2.3.
It is not clear whether the proof of Theorem 2.2 works for a discrete entry distribution $\xi$, and the proof of Theorem 2.1 of [jain2021circular] does not immediately work here on the set of compressible vectors because the structure of $\mathcal{L}$ is slightly different. This is left as an open problem. When $m_{n}$ is independent of $n$, the first statement on the least singular value of $\mathcal{L}$ was proven in a stronger form in [MR4303893] (under finite fourth moment condition) and [MR2861673] (for subGaussian entries) with a much better quantitative estimate. See also [MR2861673] for an earlier lower bound which was polynomial in $n$ for finite $m_{n}$.
3. Delocalization and rigidity estimates
Throughout the section we assume that $X$ is the general inhomogeneous matrix model in Definition 1.2.
We will use the method of free probability and concentration inequalities from [bandeira2023matrix] to derive estimates on Stieltjes transform. We first work with Gaussian entries. For any $n\times n$ matrix $A\in\mathcal{M}_{n}(\mathbb{C})$ with Gaussian entries, $A$ can be written in the form
$A:=A_{0}+\sum_{i=1}^{m}A_{i}g_{i}$
where $g_{1},\cdots,g_{m}$ are i.i.d. standard real Gaussian variables of mean 0, variance one and $A_{0},A_{1},\cdots,A_{m}$ are fixed matrices in $\mathcal{M}_{n}(\mathbb{C})$. To this model $A$ we can introduce a model in free probability
$A_{\text{free}}:=A_{0}\otimes\mathbf{1}+\sum_{i=1}^{m}A_{i}s_{i},$
where $s_{1},\cdots,s_{m}$ are a free semicircular family in some $C^{*}$ algebra $\mathcal{A}$ (see Section 4.1 of [bandeira2023matrix] for more details).
We consider the dilation matrix of $X_{z}$. Define for any $z\in\mathbb{C}$ the dilation matrix
$\mathcal{Y}_{z}:=\begin{pmatrix}0&X-zI_{n}\\ X^{*}-\bar{z}I_{n}&0\end{pmatrix}$. (3.1)
The Stieltjes transform of $\mathcal{Y}_{z}$ is defined as: for any $\eta\in\mathbb{C}_{+}:=\{z\in\mathbb{C}:\Im z>0\}$,
$\mathcal{G}_{z}(\eta):=(\mathcal{Y}_{z}-\eta I_{2n})^{-1},\quad m_{z}(\eta):=\operatorname{tr}\mathcal{G}_{z}(\eta),$
where $\operatorname{tr}$ denotes the normalized trace: for a $n\times n$ matrix $M$, $\operatorname{tr}M=\frac{1}{n}\operatorname{Tr}M$.
Our interest in the matrix $\mathcal{Y}_{z}$ stems from the fact that, a rigidity estimate on the eigenvalues of $\mathcal{Y}_{z}$ implies a rigidity estimate on singular values of $X_{z}$. We will prove a rigidity estimate for $\mathcal{Y}_{z}$ showing that not too many eigenvalues of $\mathcal{Y}_{z}$ are contained in a neighborhood of zero, which implies that except a certain amount of singular values, all other singular values of $X_{z}$ have a polynomial lower bound.
We also introduce the following free probability analogues of $\mathcal{Y}_{z}$ and $m_{z}(\eta)$. Define
$\mathcal{Y}_{z,\text{free}}:=\begin{pmatrix}0&X_{\text{free}}-zI_{n}\\ X_{\text{free}}^{*}-\bar{z}I_{n}&0\end{pmatrix}$,
$\mathcal{G}_{z,\text{free}}(\eta):=(\mathcal{Y}_{z,\text{free}}-\eta I_{2n})^{-1},\quad m_{z,\text{free}}(\eta):=\operatorname{tr}\mathcal{G}_{z,\text{free}}(\eta).$
Since the matrix $X$ has doubly stochastic variance profile, we show that $m_{z,\text{free}}(\eta)$ has a deterministic expression and is independent of the underlying graph chosen. These computations have already been done in [han2024outliers], Section 3.1 and we recall them here. By [haagerup2005new], equation 1.5 and [han2024outliers], equation 3.1, the free probability Stieltjes transform $\mathcal{G}_{z,\text{free}}(\eta)$ solves the following self consistency equation
$\mathbb{E}[\mathcal{Y}_{0}\mathcal{G}_{z,\text{free}}(\eta)\mathcal{Y}_{0}]+G_{z,\text{free}}(\eta)^{-1}+\begin{pmatrix}\eta I_{N}\quad zI_{N}\\ \bar{z}I_{N}\quad\eta I_{N}\end{pmatrix}=0.$ (3.2)
This self consistency equation uniquely characterizes $\mathcal{G}_{z,\text{free}}(\eta)$ as by [helton2007operator], Theorem 2.1, for any $\eta\in\mathbb{C}_{+}$ there is a unique solution $\mathcal{G}_{z,\text{free}}(\eta)$ to this equation such that $\mathcal{G}_{z,\text{free}}(\eta)$ has a positive imaginary part. It can be checked as in [han2024outliers] that the unique solution to (3.2) with positive imaginary part is given by
$\mathcal{G}_{z,\text{free}}(\eta)=\begin{pmatrix}a(z,\eta)I_{N}&b(z,\eta)I_{N}\\ \bar{b}(z,\eta)I_{N}&c(z,\eta)I_{N}\end{pmatrix},\quad m_{z,\text{free}}(\eta)=\frac{1}{2}(a(z,\eta)+c(z,\eta)),$ (3.3)
where $a(z,\eta)$, $b(z,\eta)$ and $c(z,\eta)$ are scalar functions depending on $z$ and $\eta$. Moreover, $a$, $b$ and $c$ solve the scalar equations (see [han2024outliers], equation (3.4))
$\begin{pmatrix}c&0\\ 0&a\end{pmatrix}+\frac{1}{ac-|b|^{2}}\begin{pmatrix}c&-b\\ -\bar{b}&a\end{pmatrix}+\begin{pmatrix}\eta&z\\ \bar{z}&\eta\end{pmatrix}=0.$ (3.4)
From this computation we see that $\mathcal{G}_{z,\text{free}}(\eta)$, and hence $m_{z,\text{free}}(\eta)$, are independent of the specific variance profile for the inhomogeneous non-Hermitian matrix $X$, and in particular the Stieltjes transform $m_{z,\text{free}}(\eta)$ for $X_{z}$ coincides with the Stieltjes transform of $G_{z}$, a shifted square $n\times n$ matrix with $i.i.d.$ Gaussian entries.
The next task is to correctly identify the rate of convergence of the Stieltjes transform $m_{z}(\eta)$ towards the deterministic limit $m_{z,\text{free}}(\eta)$.
We first assume that $X$ has independent Gaussian entries, then the following theorem directly follows from applying [bandeira2023matrix], Theorem 2.8 and Corollary 4.14.
Proposition 3.1.
Assume that $X=(x_{ij})$ is a $n\times n$ random matrix. We assume that $X$ has independent Gaussian entries and a doubly stochastic variance profile $\mathbb{E}[XX^{*}]=\mathbb{E}[X^{*}X]=\mathbf{1}$. Let
$b_{n}^{-1}:=\sup_{(i,j)\in[n]^{2}}\mathbb{E}[|x_{ij}|^{2}].$
Then with probability at least $1-n^{-100}$, we have
$\|\mathcal{G}_{z}(\eta)-\mathcal{G}_{z,\text{free}}(\eta)\|_{op}\leq\frac{C}{b_{n}|\Im\eta|^{5}}+\frac{C(\log n)^{3}}{\sqrt{b}_{n}|\Im\eta|^{2}}$ (3.5)
for some universal constant $C>0$.
In particular, for any $\mathbf{c}>0$ and for any $\Im\eta\geq(b_{n})^{-1/5}n^{\mathbf{c}}$, we have with probability at least $1-n^{-100}$,
$\|\mathcal{G}_{z}(\eta)\|_{op}\leq C_{1}$ (3.6)
where $C_{1}$ is another universal constant that can be chosen uniform for all $|z|\leq 3$ and all $|\eta|\leq 10$ (we will only consider $z$ and $C$ in this range).
Proof.
The claim (3.5) directly follows from [bandeira2023matrix], Theorem 2.8 and Corollary 4.14. The bound (3.6) is then a direct consequence of the uniform upper bound on $\mathcal{G}_{z,\text{free}}$ that can be read off from the solution of the cubic equation (3.4), see for example [girko2012theory], [MR3770875]. ∎
The upper bound on the Green function (3.6) immediately implies eigenvalue rigidity at the scale $\Im\eta\gg(b_{n})^{-1/5}$.
Corollary 3.2.
Let $X$ satisfy the assumptions in Proposition 3.1. Then for any $z\in\mathbb{C}$ and any $\mathbf{c}>0$, we can find universal constants $C_{2}$ such that with probability $1-n^{-10}$, for any interval $I\subset[-5,5]$ with length
$|I|\geq(b_{n})^{-1/5}n^{\mathbf{c}},$
we have
$\{\lambda\in I:\text{$\lambda$ is an eigenvalue of $\mathcal{Y}_{z}$}\}\leq C_{2}n|I|.$
Proof.
The proof follows from a standard argument via Helffer-Sjostrand formula, and a proof for Wigner matrices can be found in [benaych2016lectures], Section 8. We only need to make two adaptations: first, we replace Stieltjes transform of semicircle law to Stieltjes transform $m_{z,\text{free}}(\eta)$; and second, we do not take $\Im\eta\sim n^{-1+\mathbf{c}}$ but only to $\Im\eta\sim(b_{n})^{-1/5}n^{\mathbf{c}}$. Finally the constant $C_{2}$ depends on a universal upper bound on $m_{z,\text{free}}(\eta)$ over $|z|\leq 3,|\eta|\leq 10$. The details are omitted ∎
The Green function estimate (3.6) also implies eigenvector delocalization via a standard argument.
Proof of Theorem 1.9, Gaussian case.
Via a continuity argument (finding grid points and using Lipschitz continuity of $\mathcal{G}_{z}$ in $z$ and $\eta$) we can show that with probability at least $1-n^{-10}$, we have the bound (3.6) holds uniformly over all $|z|\leq 3$ and $|\eta|\leq 5$.
Consider any unit eigenvector $\psi\in\mathbb{C}^{n}$ of $X$ with eigenvalue $\lambda_{\psi}$: $(X-\lambda_{\psi})\psi=0$. Then $(0,\psi)\in\mathbb{C}^{2n}$ is an eigenvector of $Y_{\lambda_{\psi}}$ with with eigenvalue 0. By classical resolvent expansion (see for example [benaych2016lectures], Theorem 2.10), for an $n\times n$ symmetric matrix $M$ with unit eigenvectors $u_{1},\cdots,u_{n}$ associated respectively to eigenvalues $\lambda_{1},\cdots,\lambda_{n}$, denoting by $G(\eta):=(M-\eta)^{-1}$ the Green function, we have for each $i,k=1,\cdots,n$,
$\Im G_{kk}(z)=\sum_{j=1}^{n}\frac{\Im\eta}{(\lambda_{j}-\lambda_{i})^{2}+(\Im\eta)^{2}}|u_{j}(k)|^{2}\geq\frac{1}{\Im\eta}|u_{i}(k)|^{2}$
Taking $\Im\eta=(b_{n})^{-1/5}n^{\mathbf{2c}}$ for some sufficiently small $\mathbf{c}>0$, this and (3.6) shows that $\|(0,\psi)\|_{\infty}\leq(b_{n})^{-1/10}n^{\mathbf{c}}$ for each eigenvector $\psi$. Thus we have $\|\psi\|_{\infty}\leq(b_{n})^{-1/10}n^{\mathbf{c}}$.
Finally, it is easy to show (see for example [bandeira2023matrix],Theorem 2.1) that with probability $1-n^{-100}$ we have $\|X\|\leq 2.5$ (where for a matrix $X$ we use $\|X\|$ to denote its operator norm), so all eigenvalues of $X$ have modulus bounded by 3. ∎
Now we assume that $X$ has a general entry distribution. In this case we will use convergence of Stieltjes transform proved in a fairly general setting in [brailovskaya2022universality], and we need a truncation procedure to handle the tails of $\xi$.
Proposition 3.3.
Assume that $X=(x_{ij})$ is a $n\times n$ random matrix with independent entries, having mean zero and a doubly stochastic variance profile $\mathbb{E}[XX^{*}]=\mathbb{E}[X^{*}X]=\mathbf{1}$. For each $(i,j)\in[n]^{2}$ $x_{ij}$ is distributed as $b_{ij}\xi$ where $\xi$ has mean zero, variance one, finite $p$-th moment for all $p\in\mathbb{N}_{+}$. Set
$b_{n}^{-1}=\sup_{(i,j)\in[n]^{2}}b_{ij}^{2}$
and assume that $\log_{n}b_{n}>\epsilon>0$ for some $\epsilon>0$. Assume that for some (small) $\mathbf{c}^{\prime}>0$,
$|\Im\eta|^{8}b_{n}\geq n^{\mathbf{c}^{\prime}}.$
Then with probability at least $1-n^{-10}$, we have for some $\mathbf{c}>0$ depending only on $\mathbf{c}^{\prime}$ that
$\|\mathcal{G}_{z}(\eta)-\mathcal{G}_{z,\text{free}}(\eta)\|_{op}\leq\frac{n^{\mathbf{c}/2}}{b_{n}|\Im\eta|^{8}}+n^{-10},$ (3.7)
and thus we have with probability at least $1-n^{-10}$ that
$\|\mathcal{G}_{z}(\eta)\|\leq C_{1}$ (3.8)
for a universal constant $C_{1}>0$ that can be chosen uniformly for $|z|\leq 3$ and $|\eta|\leq 5$.
Proof.
We take a truncation argument. Let $\hat{\xi}:=\xi 1_{|\xi|\leq\sqrt{b_{n}}n^{-\mathbf{c}}}$ for some sufficiently small $\mathbf{c}>0$. Let $\hat{X}$ be the random matrix constructed just as $X$ but with atom distribution $\xi$ changed to $\hat{\xi}-\mathbb{E}[\hat{\xi}$]. Suppose that $\mathbb{E}[\hat{\xi}]=0$, then from the moment assumption on $\xi$, we check that with probability at least $1-n^{-50}$ we have $X=\hat{X}$ and $\mathcal{G}_{z}(\eta)=\widehat{\mathcal{G}}_{z}(\eta),$ the latter being the Stieltjes transform of the dilation (3.1) of $\widehat{X}$. Let $\widehat{\mathcal{G}}_{z,\text{free}}(\eta)$ denote the Stieltjes transform defined analogously as ${\mathcal{G}}_{z,\text{free}}(\eta)$ where we replace $X_{\text{free}}$ by $\hat{X}_{\text{free}}$: $\widehat{\mathcal{G}}_{z,\text{free}}(\eta)$ can be computed similarly via (3.2) and the only difference here is that the entries of $\hat{X}$ has mean zero and a slightly smaller variance $\mathbb{E}[|\hat{\xi}|^{2}]=1-o(1)$.
Then we first apply [brailovskaya2022universality], Theorem 2.10 to the difference $\mathbb{E}[\|\widehat{\mathcal{G}}_{z}(\eta)-\widehat{\mathcal{G}}_{z,\text{free}}(\eta)\|]$ and then apply the resolvent concentration inequality in [brailovskaya2022universality], Proposition 5.6 [as explained in [han2025circular1], Section 4, we can adapt [brailovskaya2022universality], Proposition 5.6 to be a concentration inequality for each individual entry] to bound the fluctuation of $\widehat{\mathcal{G}}_{z}(\eta)$ around its mean. Analyzing the error terms implies we should take $|\Im\eta|^{8}b_{n}\geq n^{\mathbf{c}^{\prime}}$ for some $\mathbf{c}^{\prime}>0.$ Finally we take the uniform upper bound of $\widehat{G}_{z,\text{free}}(\eta)$ for $|z|\leq 3$ and $|\eta|\leq 5$.
In general when $\mathbb{E}[\hat{\xi}]\neq 0$, by moment assumptions on $\xi$ we have that $\mathbb{E}[|\hat{\xi}|]\leq n^{-1000}$, and thus by the resolvent formula $(A-z)^{-1}-(B-z)^{-1}=(A-z)^{-1}(B-A)(B-z)^{-1}$, we can show the same estimate for $\widehat{\mathcal{G}}_{z}(\eta)$ is also the limit for $\mathcal{G}_{z}(\eta)$. ∎
Similarly to the Gaussian case, we have an immediate corollary with exactly the same proof that will be omitted:
Corollary 3.4.
Let $X$ satisfy the assumptions in Proposition 3.3. Then for any $z\in\mathbb{C}$ and any $\mathbf{c}>0$, we can find universal constants $C_{2}$ such that with probability $1-n^{-10}$, for any interval $I\subset[-5,5]$ with length
$|I|\geq(b_{n})^{-1/8}n^{\mathbf{c}},$
we have
$\{\lambda\in I:\text{$\lambda$ is an eigenvalue of $\mathcal{Y}_{z}$}\}\leq C_{2}n|I|.$
This also implies eigenvector delocalization bounds:
Proof of Theorem 1.9, non-Gaussian case.
The proof is the same as the Gaussian case, using Proposition 3.3 instead of Proposition 3.1. ∎
3.1. The Stieltjes transform convergence
We now use the already derived Green function convergence estimates in Proposition 3.1 and 3.3 to obtain the following polynomial convergence rate in Kolmogorov distance:
Theorem 3.5.
Let $X$ be defined as in Definition 1.2 but with $b_{n}\geq n^{\epsilon}$ for some small $\epsilon>0$, and with $\xi$ having mean 0, variance 1 and finite $p$-th moment for all $p$. Then for any fixed $z\in\mathbb{C}$ we can find $\zeta>0$ such that with probability 1o(1)1-o(1), we have
$\|\nu_{X_{z}}(\cdot)-\nu_{G_{z}}(\cdot)\|_{[0,\infty)}=O(n^{-\zeta}).$ (3.9)
Proof of Theorem 3.5.
Let $\widetilde{G}_{z}$ denote the dilation matrix
$\widetilde{G}_{z}:=\begin{pmatrix}0&G-zI\\ G^{*}-\bar{z}I&0\end{pmatrix}$.
Observing that $\lambda$ is an eigenvalue of $X_{z}X_{z}^{*}$ (resp. $G_{z}G_{z}^{*}$) if and only if $\pm\sqrt{\lambda}$ are both eigenvalues of $\mathcal{Y}_{z}$ (resp. $\widetilde{G}_{z}$), we see that to prove (3.9), it suffices to prove
$\|\mu_{\mathcal{Y}_{z}}(\cdot)-\mu_{\widetilde{G}_{z}}(\cdot)\|_{(-\infty,\infty)}=O(n^{-\zeta}),$ (3.10)
where we recall that $\mu_{X}$ denotes the empirical eigenvalue distribution: it differs from $\nu_{X}$ by considering the eigenvalues instead of considering squared singular values, but this change is compatible with definition of the Kolmogorov distance up to a universal constant. In other words, a bound on $\|\mu_{\mathcal{Y}_{z}}(\cdot)-\mu_{\widetilde{G}_{z}}(\cdot)\|_{(-\infty,\infty)}$ immediately implies a bound on $\|\nu_{\mathcal{Y}_{z}}(\cdot)-\nu_{\widetilde{G}_{z}}(\cdot)\|_{[0,\infty)}$, which by definition equals $\|\nu_{X_{z}}(\cdot)-\nu_{G_{z}}(\cdot)\|_{[0,\infty)}$.
The benefit of this reduction is that both $\mathcal{Y}_{z}$ and $\widetilde{G}_{z}$ have independent entries modulo symmetry. To prove (3.10), we show the Stieltjes transform of $\mathcal{Y}_{z}$ converges to the Stieltjes transform of $\widetilde{G}_{z}$ at a polynomial rate. Recall that $m_{z}(\eta):=\operatorname{tr}\mathcal{G}_{z}(\eta)$ is the Stieltjes transform of $\mathcal{Y}_{z}$, and $m_{z,\text{free}}(\eta)$ is the Stieltjes transform of $\mathcal{Y}_{z,\text{free}}$. Here the expression $m_{z,\text{free}}(\eta)$ has been explicitly computed in (3.3), (3.4).
(The Gaussian case) First assume that $X$ has Gaussian entries. If we let $g_{z,\text{free}}(\eta)$ denote the Stieltjes transform of $\widetilde{G}_{z,\text{free}}$ then we have $m_{z,\text{free}}(\eta)=g_{z,\text{free}}(\eta)$ as they both satisfy the defining relations (3.3), (3.4). Hence in the following we use $\mu_{z}(\cdot)$ to denote both $\mu_{\mathcal{Y}_{z},\text{free}}(\cdot)$ and $\mu_{\widetilde{G}_{z},\text{free}}(\cdot)$. By [MR2663633], Lemma 3.1 and Remark 3.1, $\mu_{z}(\cdot)$ has a bounded support and bounded density: for each $x\in\mathbb{R},y\geq 0:$
$|\mu_{z}((-\infty,x+y))-\mu_{z}((-\infty,x))|\leq y.$ (3.11)
Using Theorem 2.8 and Corollary 4.14 of [bandeira2023matrix], we deduce that for any $z\in\mathbb{C}$ and any $\eta$ with $\Im\eta>0$, with probability $1-n^{-100}$ we have
$|m_{z}(\eta)-m_{z,\text{free}}(\eta)|\leq\frac{(\log n)^{5}}{b_{n}|\Im\eta|^{5}}.$ (3.12)
Fix an arbitrarily large $A>0$ and Denote by $\mathcal{D}_{A}$ the region
$\mathcal{D}_{A}:=\{z\in\mathbb{C}:-A\leq\Re z\leq A,(b_{n})^{-1/6}\leq\Im z\leq 1\}.$
Using the Lipschitz continuity of $m_{z}(\eta)$ and $g_{z}(\eta)$ in $\eta$, we can upgrade the convergence in (3.12) to be uniform over $\eta\in\mathcal{D}_{A}$, with probability at least $1-n^{-10}$.
In the following we take $K>0$ sufficiently large such that $\mu_{\mathcal{Y}_{z}}(\mathbb{R}\setminus[-K,K])=0$ with probability 1o(1)1-o(1), and $\mu_{\mathcal{Y}_{z,\text{free}}}(\mathbb{R}\setminus[-K,K])=0$. Also take some $a>0$ sufficiently large. Now we use [MR2567175], Corollary B.15 to derive, for some $A>0$ large depending on $K$ and $a$, (writing $\xi=\theta+i\tau$)
$\|\mu_{\mathcal{Y}_{z}}(\cdot)-\mu_{z}(\cdot)\|_{(-\infty,\infty)}$
$\leq C\left[\int_{-A}^{A}|m_{z}(\xi)-m_{z,\text{free}}(\xi)|d\theta+\frac{1}{\tau}\sup_{x}\int_{|y|\leq 2\tau a}|\mu_{z}((-\infty,x+y])-\mu_{z}((-\infty,x])||dy\right]$
for some $C$ depending only on $A$, $a$ and $K$. Now we set $\tau=(b_{n})^{-1/6}$. To bound the first term on the second line we use (3.12), and to bound the second term on the second line use (3.11). Doing the same computation for $\widetilde{G}_{z}$ completes the proof of (3.10).
(The non-Gaussian case) For a non-Gaussian X, we can associate it with a Gaussian model having the same mean and variance, and thus $m_{z,\text{free}}(\eta)$, the associated free probability model, is uniquely determined. Then we use estimate (3.7) in place of (3.12) to complete the proof, which leads to a different value of $\zeta>0$. ∎
4. The circular law proof
In this section we collect the aforementioned results on least singular value and Stieltjes transform convergence rate, to deduce convergence of the log potential and justify the circular law. This is the technique used by [jain2021circular], and earlier versions of such argument can be found in [silverstein1995empirical], [MR2567175] and [tao2008random].
To complete the proof of circular law, we shall use the replacement principle by Tao and Vu [WOS:000281425000010].
Theorem 4.1.
Assume that for each $n$, $X$ and $G$ are two ensembles of size $n\times n$ random matrices satisfying
  1. (1)
    The quantity
    $\frac{1}{n}\|G\|_{\text{HS}}^{2}+\frac{1}{n}\|X\|_{\text{HS}}^{2}$
    is bounded in probability (resp, almost surely), where $\|\cdot\|_{\text{HS}}$ denotes the Hilbert-Schmidt norm of a square matrix.
  2. (2)
    For a.e. $z\in\mathbb{C}$ with respect to Lebesgue measure, we have
    $\frac{1}{n}\log|\det(X-zI_{n})|-\frac{1}{n}\log|\det(G-zI_{n})|$
    converges to 0 in probability (resp, almost surely).
Then $\mu_{X}-\mu_{G}$ converges in probability (resp. almost surely) to zero.
Also recall that for an $n\times n$ matrix $M$, denote by $\sigma_{1}(M)\geq\sigma_{2}(M)\cdots\geq\sigma_{n}(M)$ its singular values, then
$|\det M|=\prod_{i=1}^{n}\sigma_{i}(M).$
We will also need a useful technical lemma from [jain2021circular]:
Lemma 4.2.
Consider two probability measures $\mu,\nu$ on $\mathbb{R}$ and $0<a<b$. Then
$\left|\int_{a}^{b}\log(x)d\mu(x)-\int_{a}^{b}\log(x)d\nu(x)\right|\leq 2(|\log b|+|\log a|)\|\mu-\nu\|_{[a,b]},$
in which we define
$\|\mu-\nu\|_{[a,b]}:=\sup_{x\in[a,b]}|\mu(a,x)-\nu(a,x)|.$
Now we complete the proof of circular law in Theorem 1.5 and 1.7.
Proofof Theorem 1.5.
First consider the Gaussian case. Take $\mathbf{c}>0$ to be a sufficiently small constant to be fixed later. For any fixed $z\in\mathbb{C},z\neq 0$, we truncate the sum
$\frac{1}{n}\sum_{i=1}^{n}\log\sigma_{i}(X_{z})=\frac{1}{n}\sum_{i=1}^{n}\log\sigma_{i}(X_{z})1_{\{\sigma_{i}(X_{z})\geq(b_{n})^{-1/5}n^{\mathbf{c}}\}}+\frac{1}{n}\sum_{i=1}^{n}\log\sigma_{i}(X_{z})1_{\{\sigma_{i}(X_{z})\leq(b_{n})^{-1/5}n^{\mathbf{c}}\}}.$
By Corollary 3.2, with probability $1-n^{-100}$ there are at most $C_{2}n(b_{n})^{-1/5}n^{\mathbf{c}}$ terms in the second summation since $z\neq 0$.
Then by Theorem 2.1, with probability 1o(1)1-o(1) we have $|\log\sigma_{i}(X_{z})|\leq\frac{n^{1+3\kappa}}{b_{n}}\log b_{n}$ for each $i\in[n]$ and any small $\kappa>0$. (We take the fact that $\|X\|\leq 3$ with high probability so that $\sigma_{1}(X_{z})$ is bounded). Thus the second sum vanishes in the limit by our assumption $\log b_{n}\geq\frac{5}{6}+\epsilon$ and upon taking $\mathbf{c}>0$ and $\kappa>0$ small:
$\left|\frac{1}{n}\sum_{i=1}^{n}\log\sigma_{i}(X_{z})1_{\{\sigma_{i}(X_{z})\leq(b_{n})^{-1/5}n^{\mathbf{c}}\}}\right|\leq C_{2}\frac{1}{n}n^{\mathbf{c}}\frac{n}{(b_{n})^{1/5}}\frac{n^{1+3\kappa}}{b_{n}}=o(1),\quad n\to\infty.$ (4.1)
The same estimate in (4.1) is true if we replace $X_{z}$ by $G_{z}$ in the sum, with a much simpler proof.
Now we consider the first sum difference
$\left|\frac{1}{n}\sum_{i=1}^{n}\log\sigma_{i}(X_{z})1_{\{\sigma_{i}(X_{z})\geq(b_{n})^{-1/5}n^{\mathbf{c}}\}}-\frac{1}{n}\sum_{i=1}^{n}\log\sigma_{i}(G_{z})1_{\{\sigma_{i}(G_{z})\geq(b_{n})^{-1/5}n^{\mathbf{c}}\}}\right|$ (4.2)
$\leq 2(K_{1}+K_{2}\log n)\|\nu_{X_{z}}-\nu_{G_{z}}\|_{[0,\infty)}=o(1),\quad n\to\infty,$
where we first use Lemma 4.2 and then use Theorem 3.5, and $K_{1},K_{2}$ are two universal constants. We also used $\|X_{z}\|\leq 3+|z|$ with high probability. To sum up we have shown for any $z\in\mathbb{C}$, with probability 1o(1)1-o(1) the second criterion of Theorem 4.1 holds. The first criterion is much easier to check. Since $G$ is an i.i.d. Gaussian matrix, $\mu_{G}$ converges to the circular law, so $\mu_{X}$ also converges to circular law.
For the non-Gaussian step, we follow the same procedure but with Corollary 3.4 in place of Corollary 3.2. The only difference is now we bound, using Corollary 3.4:
$\left|\frac{1}{n}\sum_{i=1}^{n}\log\sigma_{i}(X_{z})1_{\{\sigma_{i}(X_{z})\leq(b_{n})^{-1/8}n^{\mathbf{c}}\}}\right|\leq C_{2}\frac{1}{n}n^{\mathbf{c}}\frac{n}{(b_{n})^{1/8}}\frac{n^{1+3\kappa}}{b_{n}}=o(1),\quad n\to\infty,$ (4.3)
and the quantity is vanishing to zero thanks to the assumption $\log b_{n}\geq\frac{8}{9}+\epsilon$ when $\kappa>0$ is taken very small. ∎
Proofof Theorem 1.7.
We first show that proving the circular law for $\mathcal{M}$ is equivalent to proving the circular law for $\mathcal{L}$ as defined in (2.2): this is a standard linear algebra exercise. Indeed, if $\lambda$ is an eigenvalue of $\mathcal{M}$, then $\lambda^{m_{n}}$ is an eigenvalue of $X_{1}\cdots X_{m_{n}}$ and vice versa. Meanwhile, by [MR4076784], Proposition 4.1, eigenvalues of $X_{1}\cdots X_{m_{n}}$ and eigenvalues of $\mathcal{L}^{m_{n}}$ are equal with multiplicity $m_{n}$.
Thus it suffices to prove circular law for $\mathcal{L}$. For this we follow exactly the proof of Theorem 1.5 in the non-Gaussian case where we use Theorem 2.2 for the model $\mathcal{L}$. We require that $n\geq(nm_{n})^{\frac{8}{9}+\epsilon}$, so that $m_{n}\leq n^{\frac{1}{8}-\epsilon^{\prime}}$ for some $\epsilon^{\prime}>0$. The details are omitted. ∎
5. Least singular value for linearized model
In this section, we prove Theorem 2.2.
For a sufficiently large $K>0$, denote by $\mathcal{E}_{K}$ the following event
$\mathcal{E}_{K}:=\{\forall i\in[m_{n}],\|X_{i}\|\leq K,s_{min}(X_{i})\geq n^{-5},|z|\leq K\},$
where $s_{min}(X_{i})$ is the least singular value of $X_{i}$. Then by the assumptions on $\xi$, we have, as in [jain2021circular], Lemma 2.4, that $\mathcal{E}_{K}$ holds with high probability:
Lemma 5.1.
There exists $K>0$ depending only on $\xi,z$ such that $\mathbb{P}(\mathcal{E}_{K})\geq 1-Kn^{-1}.$
For $\alpha,\beta\in(0,1)$ we define the following $z$-dependent subset of good vectors:
$L_{\alpha,\beta}:=\{v\in\mathbb{S}^{nm_{n}-1}:|\{i\in[nm_{n}]:|v_{i}|\geq\beta|\hat{z}|^{m_{n}}n^{-10m_{n}}(nm_{n})^{-1/2}\}|\geq\alpha nm_{n}\},$
where we denote by $\hat{z}:=\min(|z|,1/|z|)$ throughout this section. The exponent $|z|^{m_{n}}$ may look weird but it shows up as our matrix solves an inductive sequence with multiplications.
Denote by $\mathcal{L}_{z}=\mathcal{L}-zI_{nm_{n}}$, then we can decompose the singular value event as follows: since $s_{min}(\mathcal{L}_{z})=s_{min}(\mathcal{L}_{z}^{*})$, we switch rows and columns and write
$\mathbb{P}(\mathcal{E}_{K}\cap\{s_{min}(\mathcal{L}_{z})\leq t|z|^{m_{n}}n^{-10m_{n}}(nm_{n})^{-1/2}\})$
$\leq\mathbb{P}(\mathcal{E}_{K}\cap\{\inf_{v\in L_{\alpha,\beta}}\|\mathcal{L}_{z}^{*}v\|\leq t|z|^{m_{n}}n^{-10m_{n}}(nm_{n})^{-1/2}\})$
$+\mathbb{P}(\mathcal{E}_{K}\cap\{\inf_{v\in L_{\alpha,\beta}^{c}}\|\mathcal{L}_{z}^{*}v\|\leq t|z|^{m_{n}}n^{-10m_{n}}(nm_{n})^{-1/2}\}),$
where $L_{\alpha,\beta}^{c}$ is the complement of $L_{\alpha,\beta}$ in $\mathbb{S}^{nm_{n}-1}$. Then as in the proof of [jain2021circular], Lemma 2.5, we have the following reduction to a distance problem for vectors in $L_{\alpha,\beta}$:
Lemma 5.2.
Let $x_{i}-ze_{i}$ denote the $i$-th row of $\mathcal{L}_{z}$, and $\mathcal{H}_{i}$ denote the span of all rows except the $i$-th, for all $i\in[nm_{n}]$. Then
$\mathbb{P}(\mathcal{E}_{K}\cap\{\inf_{v\in L_{\alpha,\beta}}\|\mathcal{L}_{z}^{*}v\|\leq t|z|^{m_{n}}n^{-10m_{n}}(nm_{n})^{-1/2}\})$
$\leq\frac{1}{\alpha nm_{n}}\sum_{k=1}^{nm_{n}}\mathbb{P}(\mathcal{E}_{K}\cap\{\operatorname{dist}(x_{k}-ze_{k},\mathcal{H}_{k})\leq\beta^{-1}t\}).$
For any $v\in\mathbb{R}^{nm_{n}}$ of unit norm, we write its components as $v=(v_{[1]},v_{[2]},\cdots,v_{[m_{n}]})$ where each component $v_{[i]}\in\mathbb{R}^{n}$. Then we need the following result on the structure of normal vectors:
Proposition 5.3.
For any $z\neq 0$, on the event $\mathcal{E}_{K}$, for any vector $v\in\mathbb{S}^{nm_{n}-1}$ that is orthogonal to $\mathcal{H}_{1}$, whenever $n$ is sufficiently large (depending only on $|z|$ and $K$), we have that
$\|v_{[i]}\|\geq|\hat{z}|^{m_{n}}n^{-10m_{n}}(nm_{n})^{-1/2},\quad\forall i\in[m_{n}].$
This structural characterization is similar to [jain2021circular], Proposition 2.6 but here we crucially use the fact that $z\neq 0$ and we have only one random block per row for $\mathcal{L}_{z}$.
Proof.
The vector $v$ must satisfy that $zv_{[j]}+X_{j+1}v_{[j+1]}=0$ for each $2\leq j\leq m_{n}$. Since $v$ is a unit vector, we must be able to find some $j_{0}\in[m_{n}]$ such that $\|v_{[j_{0}]}\|\geq(m_{n})^{-1/2}$. If $j_{0}\geq 2$, using the relation $zv_{[j_{0}]}+X_{j_{0}+1}v_{[j_{0}+1]}=0$ we deduce that $\|v_{[{j_{0}+1]}}\|\geq|z|(m_{n})^{-1/2}n^{-5}$ on the event $\mathcal{E}_{K}$. Apply this relation iteratively to $j_{0}+2,j_{0}+3,\cdots$ and finally back to $\|v_{[1]}\|$ leads to the claimed bound that
$\|v_{[i]}\|\geq|z|^{m_{n}}n^{-10m_{n}}(m_{n})^{-1/2}\quad\forall j_{0}\leq i\leq m_{n}\text{ and also for }i=1,$
since we iterate for at most $m_{n}$ times in this linear system, and we cover $i=1$ by the circular update rule. Meanwhile, we consider $zv_{[j_{0}-1]}+X_{j_{0}}v_{[j_{0}]}=0$ and deduce that $\|v_{[j_{0}-1]}\|\geq|z|^{-1}(m_{n})^{-1/2}n^{-5}$. Iterating this for $j_{0}-1,j_{0}-2,\cdots,2$ verifies that
$\|v_{[i]}\|\geq|\frac{1}{z}|^{m_{n}}n^{-10m_{n}}(m_{n})^{-1/2},\quad\forall 2\leq i\leq j_{0}.$
For $j_{0}=1$ the proof is similar. ∎
Then we rule out $L_{\alpha,\beta}^{c}$ from being close to the kernel of $\mathcal{L}_{z}$ when $\alpha,\beta$ are small.
Definition 5.4.
For any $\alpha_{0},\kappa_{0}\in(0,1)$ and any integer $k\in\mathbb{N}$, we let $\operatorname{sparse}_{k}(a_{0})$ denote the set of unit vectors on $\mathbb{S}^{k-1}$ supported on at most $a_{0}k$ coordinates, and $\operatorname{Comp}(a_{0},\kappa_{0})$ denote the set of unit vectors with distance at most $\kappa_{0}$ from an $a_{0}$-sparse vector. Then denote by $\operatorname{Incomp}_{k}(a_{0},\kappa_{0})$ the complement of $\operatorname{Comp}_{k}(a_{0},\kappa_{0})$ on $\mathbb{S}^{k-1}$.
Lemma 5.5.
Let $M_{i}$ denote the $n\times 2n$ matrix given by $zI_{n},X_{i+1}]$, then we can find some $\gamma\in(0,1)$ and some $a_{0},\kappa_{0}\in(0,1)$ depending only on the entry law such that
$\mathbb{P}(\mathcal{E}_{K}\cap\{\inf_{w\in\operatorname{Comp}_{2n}(a_{0},\kappa_{0})}\|M_{i}w\|\leq\gamma\})\leq\exp(-\gamma n).$
Proof.
This argument is standard by now. We decompose $w=(w_{1},w_{2})$ into its two components, then discretize the unit sphere $\mathbb{S}^{2n-1}$, using row-wise anti-concentration of $X_{i+1}w_{2}$ to complete the proof, see for instance [jain2021circular], Lemma 2.9. ∎
Corollary 5.6.
With the given value of $\gamma$, we can find $\alpha,\beta\in(0,1)$ such that
$\mathbb{P}(\mathcal{E}_{K}\cap\{\inf_{v\in L_{\alpha,\beta}^{c}}:\|\mathcal{L}_{z}^{*}v\|\leq\gamma|\hat{z}|^{m_{n}}n^{-10m_{n}}(nm_{n})^{-1/2}\})\leq m_{n}\exp(-\gamma n).$
Proof.
On the event $\mathcal{E}_{K}$, by Proposition 5.3, for each $i$ we have
$\|(v_{[i]},v_{[i+1]})\|\geq|\hat{z}|^{m_{n}}n^{-10m_{n}}(nm_{n})^{-1/2},$
then we combine this into Lemma 5.5. To obtain a suitable value of $\alpha,\beta$, we use [rudelson2008littlewood], Lemma 3.4, which states that an incompressible vector is flat. ∎
Now we can complete the proof of Theorem 2.2.
Proof of Theorem 2.2.
We work on the event $\mathcal{E}_{K}$. Then by Corollary 5.6, it suffices to consider $\inf_{v\in L_{\alpha,\beta}}\|\mathcal{L}_{z}^{*}v\|$. Then we have, with ${x_{1}}_{[2]}$ an $n$-dimensional vector of i.i.d. entries having the law $n^{-1/2}\xi$,
$\mathbb{P}(\mathcal{E}_{K}\cap\{\operatorname{dist}(x_{1}-ze_{1},\mathcal{H}_{1})\}\leq\beta^{-1}t)\leq\sup_{w}\mathbb{P}(|\langle v_{[2]},{x_{1}}_{[2]}\rangle-w|\leq\beta^{-1}t)$
$\leq tn^{1/2}\|v_{[2]}\|^{-1}\leq tn^{1/2}|\hat{z}|^{-m_{n}}n^{10m_{n}}(nm_{n})^{1/2},$
where the second inequality follows from the density assumption on $\xi$ and Theorem 1.1 of [rudelson2015small]. This completes the proof. ∎